### v1

# A GNN-based Spectral Filtering Mechanism for Imbalance Classification in Network Digital Twin

Abstract
Graph Neural Networks are gaining attention in Fifth-Generation (5G) core network digital twins, which are data-driven complex systems with numerous components. Analyzing these data can be challenging due to rare failure types, leading to imbalanced classification in multiclass settings. Digital twins of 5G networks increasingly employ graph classification as the main method for identifying failure types. However, the skewed distribution of failure occurrences is a major class imbalance issue that prevents effective graph data mining. Previous studies have not sufficiently tackled this complex problem. In this paper, we propose Class-Fourier Graph Neural Network (CF-GNN) introduces a class-oriented spectral filtering mechanism that ensures precise classification by estimating a unique spectral filter for each class. We employ eigenvalue and eigenvector spectral filtering to capture and adapt to variations in the minority classes, ensuring accurate class-specific feature discrimination, and adept at graph representation learning for complex local structures among neighbors in an end-to-end setting. Extensive experiments have demonstrated that the proposed CF-GNN could help with both the creation of new techniques for enhancing classifiers and the investigation of the characteristics of the multi-class imbalanced data in network digital twin system.
keywords:
Graph neural network , Multi-class imbalanced , Eigenvalue , Graph representation learning , Network Digital Twin
journal: Internet of Things
[…]
[1]organization=Department of Intelligent Electronics and Computer Engineering, Chonnam National University,addressline=Buk-gu, postcode=61186, state=Gwangju, country=South Korea
[…]
[3]organization=Department of Computer Science, Ahmadu Bello University,postcode=810241, state=Zaria, country=Nigeria
[…]
[4]organization=Department of Computational and Data Science, Astana IT University, Astana, Kazakhstan
{graphicalabstract}[Uncaptioned image]
{highlights}
We proposed a novel CF-GNN to address class imbalance tasks in multiclass 5G network digital twins.
CF-GNN employs the concepts of eigenvalues and eigenvectors to identify minority classes in a multiclass failure state.
Incorporating spectral filtering influenced by the global graph structure and local neighborhood information during the filtering process.
The proposed model showcased the robust performance of the network digital twin-unbalanced datasets.
1 Introduction
The emergence of fifth-generation (5G) networks, and robust core network systems are paramount for the evolution to Industry 5.0. According to recent studies, 5G connections are projected to increase over 100-fold, from approximately 13 million in 2018 to 2.6 billion by 2025 [1]. This growth is expected to continue with sixth-generation (6G) networks and the advent of the Industry 5.0 era, characterized by the vision of the Internet of Everything. In this context, network failure classification for 5G and beyond is essential to optime the network performance. As the backbone of communication infrastructure, reliable and resilient networks are essential for seamless operation in diverse applications in environments ranging from daily life to industrial settings. The core network management center can identify and classify network failures by applying deep learning algorithms and advanced analytics, enabling preemptive measures to ensure uninterrupted service delivery [2].
The digital twin (DT) is a powerful technology that has emerged as a promising solution for connecting physical spaces with digital systems [3, 4]. This technology uses historical data and real-time operational information to create digital replicas of physical entities, such as devices, machines, and objects. Research on DTs [5] and [6] has garnered significant attention, with numerous researchers exploring its applications in such fields as aviation, 6G networks, intelligent manufacturing, and data generation [7, 8]. Additionally, DT has been recognized as one of the 12 representative use cases for future networks by the International Telecommunication Union Network 2030 focus group [9]. However, setting up a replica of a network event to mimic every potential pattern in network DTs is challenging.
Numerous tasks related to graphs classification have been extensively studied, including node classification [10], graph classification [11], and link predictions [12]. Moreover, a variety of advanced GNN methodologies, such as graph pooling [13], graph convolution [14], and graph attention mechanisms [15], have been suggested to enhance the performance of GNNs. Despite these advancements, there remains a scarcity of graph models for communication network domains [16], which is crucial for understanding GNNs. Several methods, such as RouteNet [17] and RouteNet-Fermi [18], have been introduced to model the NDT for network performance. The benefit of data-driven models [19] is their ability to accurately capture a wide range of complex network nodes with unprecedented precision, as they are trained on real-world data [20]. Nevertheless, none of these recent studies have focused on data-driven network digital twin approach to tackle the imbalance node classification tasks in the NDT domain.
Class-Imbalanced across many application fields, including the Digital Twin [21], Fraud Detection [22], for computing [23], and recently the synthetic oversampling [24]. GNN models that are robust to class imbalance must be developed to avoid biases towards majority classes while maintaining the ability to generalize over minority classes. Traditional techniques for addressing the class imbalance, including oversampling [25], and sampling the classifiers [26], have limitations when applied to graph-structured data because they do not account for the inherent graph structure. Thus, many scenarios have been developed to tackle class imbalance within semi-supervised node classification. Node classification with synthetic oversampling methods, GraphSMOTE [27, 28], and GATE-GNN [29] generates new nodes and employ a variety of methods to connect them to the existing graph. Furthermore, these methods often fail to fully utilize the untapped potential of the abundant labeled and unlabeled nodes in the graph, which limits their effectiveness in real-world applications.
In this study, we present a novel Class-Fourier Graph Neural Network filtering (CF-GNN) framework designed to address the challenges posed by Imbalanced node classification tasks. Our proposed method leverage node-specific spectral filtering to ensure that minority-class nodes receive attention without being overtaken by majority-class nodes. By enabling the filter to focus on particular nodes. This helps to emphasize overshadowed classes, which is particular beneficial for imbalanced datasets. the localized filtering concentrates on the distinct features of nodes. This ensures that class-specific changes can be applied to nodes that belong to minority classes, enhancing their representation in training and prediction. Additionally, the filtering process, being aware of class-specific characteristics in the spectral domain, reduces the likelihood of minority-class nodes being overwhelmed by dominant classes. During training, the model employs class-weighted loss functions, which penalize the misclassification of minority-class samples more heavily. This forces the network to learn better representations of underrepresented classes. The spectral filtering procedure of the CF-GNN explicitly integrates a multiclass structure, transforming features in a class-aware way by the use of spectral coefficients and eigenvalue-based distinctions between classes. Our experimental evaluations on two 5G NDT datasets demonstrate how well the proposed CF-GNN framework compared to state-of-the-art GNN techniques designed to mitigate class imbalance problems.
The contributions of this work and the proposed CF-GNN include the following:
  • 1.
    The proposed CF-GNN employs a spectral attention technique in the Fourier domain to extract complex structural components that conventional spatial-based GNN may overlook.
  • 2.
    CF-GNN utilizes class-specific spectral filtering to enhance classification accuracy by highlighting significant feature components associated with certain classes. By iteratively improving the eigenvector estimates, this procedure ensure accurate spectral analysis of the graph structure.
  • 3.
    Introduces the theoretical Twin-GFT for multidimensional spectral representation, enabling scalable failure detection and improving prediction robustness through ensemble classification.
1.1 Related work
Graph learning-driven failure predictions have been gaining attention due to their effectiveness in tackling the massive volumes of data generated by 5G networks and their ability to understand the complex structural characteristics of these networks. This section explores three areas of research: one focusing on learning-based 5G failure prediction, imbalanced learning techniques, and Graph Fourier transform (GFT) to capture the complex structure of NDTs.
1.1.1 Learning-based failure prediction
The primary advantage of applying deep graph learning to network modeling [30] lies in its data-driven approach, which enable it to capture the complex nature of the real-world networks. Most existing research in [31] and [32] uses fully connected conventional neural networks. However, the primary constraint of these methods is their non-generalization to alternative network topologies and configurations, such as routing. In this context, more recent studies have proposed advanced neural network models, such as GNNs [33, 34], convolutional neural networks [32], and variational autoencoders [35] for robustness in imbalanced datasets common to cybersecurity domains. Despite their advancements, these models have different objectives and overlook critical aspects of real-world networks from the twin perspective. The concept of broadly applying neural networks to graphs has gained significant attention. For example, convolutional networks have been extended to graphs in the spectral domain [36], where filters are applied to frequency modes using GFT. The eigenvector matrix of a graph Laplacian must be multiplied to achieve this transformation. One study [37] parameterized the spectrum filters as Chebyshev polynomials of eigenvalues, resulting in efficient and localized filters and reducing the computational burden. However, a drawback of these spectral formulations are disadvantageous is that they are limited to graphs with a single structure since they depend on the fixed spectrum of the graph Laplacian. In contrast, spatial formulations are not constrained by a specific graph topology. Generalizing neural networks to graphs is studied in two ways: a) given a single graph structure, or labels of individual nodes [38, 36, 39, 40, 37, 41] b) given a set of graphs with different structure and sizes, the goal is learn predictions of the class labels of the graphs are learned [42, 43, 44]. Inspired by the ITU Challenge ”ITU-ML5G-PS-008: Network Failure Classification Model Using the Digital Twin Network” [45], our research aims to leverage GNNs, enhanced with the CF-GNN, to analyze interconnected systems and networks in an end-to-end manner. By integrating class-oriented spectral filtering, the proposed CF-GNN framework enables precise classification of network failure types and accurate identification of failure points. The approach improves accuracy and adaptability compared to conventional deep learning models, effectively capturing both global consistency and localized variations within complex 5G network structures.
1.1.2 Imbalance learning technique
Data-level and algorithm-level methods are the two main categories of class-imbalanced learning techniques. Before building classifiers, data-level approaches preprocess training data to reduce inequality [46]. These strategies include undersampling majority classes and oversampling minority classes. On the other hand, algorithm-level approaches address the issue of class imbalance by modifying the model’s fundamental learning decision-making process. Algorithm-level techniques can be broadly classified into three categories: threshold moving, cost-sensitive learning, and new loss functions [47]. Several approaches have been proposed at the data level. SMOTE [25] is one such technique that generates synthetic minority samples in the feature space by interpolating existing minority samples and their nearest minority neighbors. However, one of the primary limitations of SMOTE is that it creates new synthetic samples without considering the neighborhood samples of the majority classes, which can lead to increased class overlapping and extra noise [48]. Based on the SMOTE principle, many variations have been proposed, such as Borderline-SMOTE [49] and Safe-Level-SMOTE [50], which enhance the original approach by taking majority class neighbors into account. At the same time, Safe-Level-SMOTE creates safe zones to prevent oversampling in overlapping or noisy regions, whereas Borderline-SMOTE only samples the minority samples close to the class borders. To address class-imbalanced challenges, [51] modified the parameters of the model learning process that favored classes with fewer samples. For example, Focal Loss, introduced in [52], allows minority samples to contribute more to the loss function. Additionally, a novel loss function called Mean Squared False Error (MSFE) was proposed in [53] to train deep networks on imbalanced datasets.
This study enables us to handle the class-imbalanced problem on graph data, two new models have been developed recently: We employ some of the settings in GATE-GNN [29] dynamically modifies the weights of various GNN modules and Reweight [54] that focuses on class-incremental learning on dynamic real-world environments on the dual imbalances.
The paper was structured as follows: The notation and preliminaries are introduced in Section (2). In Section (3), we investigate the graph filtering mechanism of the proposed method and the 5G network digital twin datasets. The experiment and evaluation are provided in Section (4). details the datasets and results. We finally concluded in Section (5).
2 Notations and Preliminaries
We formulated the network failure injection as a classification problem, focusing on learning meaningful embeddings from the generated datasets to increase detection accuracy and identify different failure types. In the network structure, we define a graph $G=(V,E)$ with an adjacent matrix $A=\{e_{vu}|\forall v,u\in V\}\in\mathbb{R}^{N\times N}$, where $N$ represents the number of nodes, letting $G=(V,E)$ represent an input graph with a set of nodes $V$. If $e_{vu}=1$, then $(v,u)\in E$; otherwise, $(v,u)\in E^{\prime}$. We assume the set of feature vectors associated with each network component $v$ are $D$-dimensional represented by $F=\{f_{v}|v\in V\}\in\mathbb{R}^{D}$. We considered the failure points as nodes in the network graph for network failure classification. Each node is associated with various features, such as performance metrics or operational data, reflecting its behavior within the network. The edges connecting pairs of nodes represent the dependencies and relationships between them.
Figure 1: The architectural framework of the proposed technique for imbalanced classification of 5G network of digital twin
The proposed framework in Figure 1 illustrates the proposed model. We applied CF-GNN to examine individual nodes with different types of features (e.g., phys-address, oper-status, etc.). The GNN takes the provided NDT data as input. The GNN retrieves fine-grained information about node interactions and local graph structures, while the Class-Fourier Transform (CFT) efficiently denoizes features while capturing global network attributes. The graph module then processes these inferred relationships to create a multi-layer graph representation. The feature extraction module applies Fourier transform algorithm to convert 5G network data into useful information. To achieve accurate multi-class classification results, the proposed CF-GNN integrate the extracted features with the learned graph structure.
Table 1: Explanation of Graph and Laplacian Variables
Variable Explanation
$G=(V,E)$ Graph representing the NDT system
$V_{n}$ Set of nodes in graph $G_{n}$
$N$ Total number of nodes
$N_{n}$ Number of nodes in $G_{n}$
$L_{n}$ Laplacian matrix of $G_{n}$
$\lambda^{(n)}_{k}$ $k$-th eigenvalue of $\mathcal{L}^{(n)}$
$\phi^{(n)}_{k}$ Eigenfunctions of the Laplacian
$\mathcal{G}_{1}\times\mathcal{G}_{2}$ Product graph with node features
$\sigma(L_{1}\oplus L_{2})$ Eigenvalues of $L_{1}\oplus L_{2}$
$(i_{1},i_{2})$ Node indices in $\mathcal{G}_{1}$ and $\mathcal{G}_{2}$
$(k_{1},k_{2})$ Indices for eigenvalues and eigenfunctions
$y_{ij}$ One-hot label component for node $v_{i}$
$f_{ij}$ Predicted probability for node $v_{i}$ in class $j$
$W_{fc1},b_{fc1}$ Parameters of the GFT linear layer
$W_{mp1},b_{mp1}$ Parameters of the message passing layer
$h_{i}$ Hidden representation at layer $i$
$\hat{m},\hat{v}$ Adam optimizer moment estimates
$\eta$ Learning rate
$y$ Output vector of size $(N,C)$
$\theta$ Parameters of the GNN model
$f^{(l)}$ Aggregation function at layer $l$
$N(v_{i})$ Neighbors of node $v_{i}$
$h_{i}^{(0)}=x_{i}$ Initial feature vector for node $v_{i}$
2.1 Graph Neural Network
The GNN can transform the graph structure data into standard representations, making them suitable for input into neural networks for training. We employ the concept of class spectral filtering introduced in [55] and [56] to enhance the learning process. This approach allows the GNN to efficiently propagate the information from the nodes and edges to their neighboring nodes. To achieve this, we use the neighbor aggregation approach in [57]. The network infrastructure is modeled as a graph $G=(V,E)$, where $V$ represents the set of nodes (network components) and $E$ represents the set of edges (connections). Each node $v_{i}$ in the graph is associated with a feature vector $x_{i}$, which captures relevant information about the corresponding network component. The graph can be represented using an adjacency matrix $A$, where $A_{ij}=1$ if a connection exists between nodes $v_{i}$ and $v_{j}$, and $A_{ij}=0$ otherwise. Node features are stored in the feature matrix $X$, where $X$ has the dimensions $|V|\times D$, with $|V|$ represents the number of nodes, and $D$ denotes the dimensionality of the feature vectors $X=[x_{1},x_{2},\ldots,x_{|V|}]$.
Many network components depend upon one another due to their interconnectivity [17]. The GNN architecture consists of multiple layers, each processing information from neighboring vertices to extract hierarchical representations. We let $h_{i}^{(l)}$ denote the representation of vertex $v_{i}$ at layer $l$ of the GNN. Information is aggregated from the neighboring vertices and transformed using learnable parameters. The update equation for $h_{i}^{(l)}$ can be expressed as follow:
$h_{i}^{(l)}=f^{(l)}\left(h_{i}^{(l-1)},\{h_{j}^{(l-1)}\}_{j\in N(v_{i})}\right),$ (1)
where $N(v_{i})$ represents the set of neighboring vertices of $v_{i}$.
One of the biggest challenges is creating network digital twin datasets that accurately replicate the real-world failure scenario. This difficulty arises to several factors, such as the limited availability of real-world system data and the inherent complexity of simulating diverse failure types.
2.2 Graph Fourier Transform
We considered an undirected weighted graph $G=(V,E,w)$, where $V=\{0,1,\ldots,N-1\}$ is the vertex set, $E$ is the set of edges, and $w(i,j)$ is the weight function satisfying $w(i,j)=0$ for any $(i,j)$notin $E$. We assume all graphs are simple, meaning they have no loops or multiple edges.
Three matrices associated with $G$ are important: the adjacency matrix $W$, a degree matrix $D$, and a Laplacian matrix $L=D-W$. The Laplacian matrix is essential for Graph Fourier Transforms (GFTs).
The Laplacian matrix $L$ is real, symmetric, and positive-semidefinite, thus, it possesses nonnegative eigenvalues $\lambda_{0},\ldots,\lambda_{N-1}$ and the corresponding orthonormal eigenfunctions $u_{0},\ldots,u_{N-1}$. These eigenfunctions satisfy the following equation:
$L\begin{bmatrix}u_{k}(0)\\ \vdots\\ u_{k}(N-1)\end{bmatrix}=\lambda_{k}\begin{bmatrix}u_{k}(0)\\ \vdots\\ u_{k}(N-1)\end{bmatrix}$ (2)
For $k=0,\ldots,N-1$, orthonormality implies that the sum of products of corresponding eigenfunctions is equal to the Kronecker delta function $\delta(i,j)$, where $\delta(i,j)$ equals 1 if $i=j$ and is zero otherwise. This study assumes that eigenvalues are arranged in ascending order, such that $\lambda_{0}\leq\ldots\leq\lambda_{N-1}$. In addition, $\lambda_{0}$ is strictly zero because the sum of rows in $L$ equals zero. The spectrum of the matrix, denoted as $\{\lambda_{k}\}_{k=0}^{N-1}$, is represented as $\sigma(L)$. The GFT in [58] of a graph feature $f:V\rightarrow\mathbb{R}$ is defined as $\hat{f}:\sigma(L)\rightarrow\mathbb{C}$, where $\sigma(L)$ represents the spectrum of the graph Laplacian $L$. It is expressed as follows:
$\hat{f}(\lambda_{k})=\langle f,u_{k}\rangle=\sum_{i=0}^{N-1}f(i)u_{k}(i),$ (3)
for $k=0,\ldots,N-1$. where, $u_{k}$ denotes the orthonormal eigenfunctions of $L$, and the GFT represents a feature expansion using these eigenfunctions. The inverse GFT is given by,
$f(i)=\sum_{k=0}^{N-1}\hat{f}(\lambda_{k})u_{k}(i),$ (4)
which reconstruct the original function $f(i)$ from its spectral components. Interestingly, the GFT’s basic function in spectral graph research is highlighted by the fact that it is equal to the Discrete Fourier Transform (DFT) on cycle graphs.
When a graph Laplacian has non-distinct eigenvalues, the functions generated by the GFT may not be well-defined, resulting in multi-valued functions. For instance, if two orthonormal eigenfunctions, $u$ and $u_{0}$ correspond to the same eigenvalue $\lambda$, then the spectral component $\hat{f}(\lambda)$ can have two distinct values: $\langle f,u\rangle$ and $\langle f,u_{0}\rangle$.
2.3 Digital Twin Spectral Filtering
The Cartesian product $G_{1}\square G_{2}$ of graphs $G_{1}=(V_{1},E_{1},w_{1})$ and $G_{2}=(V_{2},E_{2},w_{2})$ is a graph with the vertex set $V_{1}\times V_{2}$ and the edge set $E$ defined as follows. For any $(i_{1},i_{2})$ and $(j_{1},j_{2})$ in the vertex set $V_{1}\times V_{2}$, these are connected by an edge if either $(i_{1},j_{1})$ is in $E_{1}$ and $i_{2}=j_{2}$ or $i_{1}=j_{1}$ and $(i_{2},j_{2})$ is in $E_{2}$. The weight function $w$ is defined as follows:
$w((i_{1},i_{2}),(j_{1},j_{2}))=w_{1}(i_{1},j_{1})\delta(i_{2},j_{2})+\delta(i_{1},j_{1})w_{2}(i_{2},j_{2})$ (5)
where, $\delta$ denotes the Kronecker delta function. The graphs $G_{1}$ and $G_{2}$ are referred to as the factor graphs of $G_{1}\square G_{2}$.
It has been established in [59] that the adjacency, degree, and Laplacian matrices of Cartesian product graph can be derived form those of its factor graphs. Two factor graphs, $G_{1}$ and $G_{2}$, with vertex sets $V_{1}=\{0,1,\ldots,N_{1}-1\}$ and $V_{2}=\{0,1,\ldots,N_{2}-1\}$ respectively are considered. Each with adjacency matrix $W_{1}$ and $W_{2}$, degree matrix $D_{1}$ and $D_{2}$, and Laplacian matrix $L_{1}$ and $L_{2}$. When the vertices of the Cartesian product graph are ordered lexicographically, such as $(0,0),(0,1),(0,2),\ldots,(N_{1}-1,N_{2}-1)$, the adjacency, degree, and Laplacian matrices of $G_{1}\square G_{2}$ can be expressed as $W_{1}\oplus W_{2}$, $D_{1}\oplus D_{2}$, and $L_{1}\oplus L_{2}$, respectively, where operator $\oplus$ denotes the Kronecker sum.
Definition 1: The Kronecker sum is defined by $A\oplus B=A\otimes I_{n}+I_{m}\otimes B$ for matrices $A\in\mathbb{R}^{m\times m}$ and $B\in\mathbb{R}^{n\times n}$, where $I_{n}$ represents the identity matrix of size $n$.
The Kronecker sum has a valuable characteristic that decomposes an eigenproblem involving the Laplacian matrix of a product graph into eigenproblems of Laplacian matrices of the factor graphs. We assumed that the Laplacian matrix $L_{n}$ of each factor graph has nonnegative eigenvalues $\{\lambda^{(n)}_{k}\}_{k=0}^{N_{n}-1}$ and orthonormal eigenfunctions $\{u^{(n)}_{k}\}_{k=0}^{N_{n}-1}$ for $n=1,2$. In this case, the Kronecker sum $L_{1}\oplus L_{2}$ yields an eigenvalue of $\lambda^{(1)}_{k}+\lambda^{(2)}_{k}$ and the corresponding eigenfunction $u^{(1)}_{k}\otimes u^{(2)}_{k}:V_{1}\times V_{2}\rightarrow\mathbb{C}$, where $\oplus$ denotes the Kronecker sum. This eigenfunction satisfies the following:
$(L_{1}\oplus L_{2})\left[\begin{array}[]{c}u^{(1)}_{k}(0)u^{(2)}_{k}(0)\\ u^{(1)}_{k}(0)u^{(2)}_{k}(1)\\ \vdots\\ u^{(1)}_{k}(N_{1}-1)u^{(2)}_{k}(N_{2}-1)\end{array}\right]$ (10)
$=(\lambda^{(1)}_{k}+\lambda^{(2)}_{k})\left[\begin{array}[]{c}u^{(1)}_{k}(0)u^{(2)}_{k}(0)\\ u^{(1)}_{k}(0)u^{(2)}_{k}(1)\\ \vdots\\ u^{(1)}_{k}(N_{1}-1)u^{(2)}_{k}(N_{2}-1)\end{array}\right]$ (15)
This property holds for any $k_{1}=0,\ldots,N_{1}-1$ and $k_{2}=0,\ldots,N_{2}-1$. The resulting eigenvalues $\{\lambda^{(1)}_{k}+\lambda^{(2)}_{k}\}$ are nonnegative, and the corresponding eigenfunctions $\{u^{(1)}_{k}\otimes u^{(2)}_{k}\}$ are orthonormal. These conclusions can be straightforwardly derived from the fundamental properties of the Kronecker product.
A similar approach applies to decomposing an eigenproblem related to an adjacency matrix of a product graph. If the adjacency matrix $W_{n}$ has eigenvalues $\{\mu^{(n)}_{k}\}_{k=0}^{N_{n}-1}$ and orthonormal eigenfunctions $\{\mathbf{v}^{(n)}_{k}\}_{k=0}^{N_{n}-1}$ for $n=1,2$, the Kronecker sum $W_{1}\oplus W_{2}$ yields the eigenvalues $\{\mu^{(1)}_{k}+\mu^{(2)}_{l}\}$ and corresponding eigenfunctions $\{\mathbf{v}^{(1)}_{k}\otimes\mathbf{v}^{(2)}_{l}\}$ defined on $V_{1}\times V_{2}$.
The Cartesian product graph is represented as $G_{1}\times G_{2}$, where $G_{n}$ for $n=1,2$ is an undirected weighted graph with a vertex set $V_{n}=\{0,1,\dots,N_{n}-1\}$. Assuming its graph Laplacian $L_{n}$ has eigenvalues $\{\lambda_{k}^{(n)}\}_{k=0}^{N_{n}-1}$ and corresponding orthonormal eigenfunctions $\{u_{k}^{(n)}\}_{k=0}^{N_{n}-1}$. The GFT of a graph feature $f:V_{1}\times V_{2}\rightarrow\mathbb{R}$ on the product graph $G_{1}\times G_{2}$ is represented by:
$\hat{f}:(\sigma(L_{1})\otimes\sigma(L_{2}))\rightarrow\mathbb{C}$ (16)
$\hat{f}(\lambda^{(1)}_{k_{1}}+\lambda^{(2)}_{k_{2}})=\sum_{i_{1}=0}^{N_{1}-1}\sum_{i_{2}=0}^{N_{2}-1}f(i_{1},i_{2})\cdot u^{(1)}_{k_{1}}(i_{1})\cdot u^{(2)}_{k_{2}}(i_{2})$ (17)
for $k_{1}=0,\dots,N_{1}-1$ and $k_{2}=0,\dots,N_{2}-1$, and its inverse is:
$f(i_{1},i_{2})=\sum_{k_{1}=0}^{N_{1}-1}\sum_{k_{2}=0}^{N_{2}-1}\hat{f}(\lambda^{(1)}_{k_{1}}+\lambda^{(2)}_{k_{2}})\cdot u^{(1)}_{k_{1}}(i_{1})\cdot u^{(2)}_{k_{2}}(i_{2})$ (18)
Thus, we interpret features on a Cartesian product graph as 2D features and propose a Twin GFT, enabling multiclass failure classification in Network Digital Twin environments.
Definition 1 The Twin Graph Fourier Transform (Twin-GFT) of a function $f:V_{1}\times V_{2}\rightarrow\mathbb{R}$ on a Cartesian product graph $G_{1}\square G_{2}$ is a spectral representation $\hat{f}:\sigma(L_{1})\times\sigma(L_{2})\rightarrow\mathbb{C}$ defined as follows:
$\hat{f}(\lambda_{11},\lambda_{22})=\sum_{i_{1}=0}^{N_{1}-1}\sum_{i_{2}=0}^{N_{2}-1}f(i_{1},i_{2})\cdot u_{k_{11}}^{(1)}(i_{1})\cdot u_{k_{22}}^{(2)}(i_{2})$ (19)
for $k_{11}=0,\dots,N_{1}-1$ and $k_{22}=0,\dots,N_{2}-1$, and its inverse is given by:
$f(i_{1},i_{2})=\sum_{k_{1}=0}^{N_{1}-1}\sum_{k_{2}=0}^{N_{2}-1}\hat{f}(\lambda_{k_{11}}^{(1)}+\lambda_{k_{22}}^{(2)})\cdot u_{k_{11}}^{(1)}(i_{1})\cdot u_{k_{22}}^{(2)}(i_{2})$ (20)
for $i_{1}=0,\dots,N_{1}-1$ and $i_{2}=0,\dots,N_{2}-1$.
The twin-GFT can be represented as a series of matrix-matrix multiplications. Using $N_{1}\times N_{2}$ matrices $F=(f(i_{1},i_{2}))_{i_{1},i_{2}}$ and $\hat{F}=(\hat{f}(\lambda_{k_{11}}^{(1)}+\lambda_{k_{22}}^{(2)}))_{k_{11},k_{22}}$, the twin-GFT applied to the features $f$ is expressed as follows:
$\hat{F}=U_{1}^{*}FU_{2},$ (21)
where $U_{n}$ denotes an $N_{n}\times N_{n}$ unitary matrix with $(i,k)$-th element $u_{k}^{(n)}(i)$ for $n=1,2$. Then its inverse transform is given by:
$F=U_{1}\hat{F}^{2}U_{2}^{*},$ (22)
The twin GFT has connections to existing transformations. First, when both factor graphs are cycle graphs, the twin GFT can be equivalent to the 2D-Discrete Fourier Transform and the 2D-GFT. Second, when a factor graph is a cycle graph, some cycles might be nested within other cycles, creating a complex network structure. The twin GFT is known as a joint graph and temporal Fourier transform which generalizes these existing transformations.
3 PROPOSED CF-GNN
Class-Fourier graph Neural Networks (CF-GNNs) perform well with both homophilic and heterophilic graphs and have shown good expressive capacities. However, current methods often lack adaptability. In particular, conventional approaches use a single, globally shared spectral filter $\hat{g}$, trained across the entire graph, with fixed frequency coefficients $\{\gamma_{k}\}_{k=0}^{K}$. Node-specific changes are not considered in the global filter, which is unable to distinguish between the distinct local structures connected to each node during the filtering process. Although polynomial-parameterized spectral filter learning offers a certain level of localization, it might not adequately capture subtle and diverse local structural patterns. It is intuitively reasonable to train a class-specific filter $\hat{g}_{i}(\lambda_{l})$ rather than depending on a globally consistent spectral filter $\hat{g}(\lambda_{l})$. A more efficient method is provided by $\hat{g}_{i}(\lambda_{l})$, tailored for each node $i$ to adaptively capture its local patterns. To mitigate this limitation, we revisit globally consistent spectral graph transforms in this section and propose a spectral filter learning strategy in the context of CF-GNNs.
3.1 Class-Fourier Transform Filtering
We present an adaptive localized spectral filtering method designed for Class-Fourier Transform Filtering on graph $G$, which the generalized translation operator inspires in Graph Signal Processing (GSP) in [60]. By leveraging polynomial-parameterized spectral filtering from [61], this technique ensures a more accurate and flexible representation of local patterns by fully accounting for the unique influence of the node at which the filter is applied. We introduce a generalized translation operator $T_{i}$ for any signal $g\in\mathbb{R}^{n}$ defined on a given graph $G$ and any $i\in\{0,1,\dots,n-1\}$: $T_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$
Using a Kronecker delta function in (17) and (18) $\delta_{i}$ centered at the $i$-th node $v_{i}$, we express generalized convolution as:
$T_{i}(g):=\sqrt{N}\,(g\ast\delta_{i})=\sqrt{N}\sum_{l=1}^{n}u_{l}u_{l}^{\top}(i)\hat{g}(\lambda_{l})$ (23)
where $u_{l}^{\top}(i)$ denotes the $i$-th element of the eigenvector $u_{l}$ corresponding to eigenvalue $\lambda_{l}$, and $\hat{g}(\lambda_{l})$ represents the spectral domain representation of the signal $g$.
In the context of Class-Fourier Transform Filtering, adaptive local filtering is achieved by first centering the filter signal $g$ at the target node $v_{i}$ using the generalized translation operator $T_{i}$, and then performing spectral convolution with $x$ as follows:
$x\ast GT_{i}(g)=\sum_{l=1}^{N}u_{l}\,x^{(\lambda_{l})}u_{l}^{\top}(i)g^{(\lambda_{l})}$ (24)
where $x^{(\lambda_{l})}$ is the spectral domain representation of the input signal $x$, and $u_{l}^{\top}(i)$ is the $i$-th element of the eigenvector $u_{l}$, which corresponds to the eigenvalue $\lambda_{l}$.
Establishing that: $\hat{g}_{i}(\lambda_{l})=\sqrt{N}\,u_{l}(i)\,\hat{g}(\lambda_{l})$ in 24, the equation above becomes:
$x\ast GT_{i}(g)=\sum_{l=1}^{N}u_{l}\,x^{(\lambda_{l})}g_{i}^{(\lambda_{l})}$ (25)
As evident from the definition of $\hat{g}_{i}(\lambda_{l})$, $\hat{g}_{i}$ is a function associated with $U$. Thus, by approximating $U$, we can further establish the relationship between $\hat{g}_{i}$ and $x$. Recall the Fourier transform of an inverted graph, which allows us to deduce that $x_{i}=U_{i:}\cdot\hat{x},$ where $U_{i:}$ denotes $U$’s $i$-th row. Therefore, $U_{i:}$ can be approximately from $x_{i}$ using $U_{i:}\approx x_{i}q,$where $\hat{x}$’s pseudoinverse is $q=\text{pinv}(\hat{x})\in\mathbb{R}^{n}$. Meanwhile, take notice that $u_{l}(i)=\tilde{u}_{l}(i)$, which implies that $\hat{g}_{i}(\lambda_{l})$ can be approximately expressed as $\hat{g}_{i}(\lambda_{l})\approx\ldots$.
$\hat{g}_{i}(\lambda_{l})\approx\tilde{g}_{i}(\lambda_{l})=\sqrt{N}(x_{i}q_{l})\hat{g}(\lambda_{l})$ (26)
Where the $q_{l}$ denotes the $l$-th element of $\mathbf{q}$. Consequently, a particular filter $\tilde{g}_{i}$ corresponding to node $v_{i}$, can approximate $\hat{g}_{i}$ while simultaneously incorporating the influence of feature $x_{i}$ associated with to node $v_{i}$ in the filter coefficient estimation process.
Without loss of generality, the $K$-order polynomial approximation $\hat{P}_{K}(\cdot)$ can be utilized to directly parameterize $\tilde{g}_{i}$ without sacrificing generality; that is,$\tilde{g}_{i}=\sum_{k=0}^{K}\eta_{i,k}\hat{p}_{k}(\Lambda),$ where $\eta_{i,k}$ represents the coefficient of $\hat{p}_{k}(\Lambda)$ for the $\tilde{g}_{i}$ filter. We only pay attention to the convolution value of the filter placed at each node is our focus, i.e.,$z_{i}=\delta_{i}(U\tilde{g}_{i}U^{\top}x).$ In this case, we utilize $\delta_{i}$ as an indicator vector $\delta_{i}=[0,\ldots,1_{i},\ldots,0]\in\mathbb{R}^{n},$ which represents a row vector with only the $i$-th element being 1 and the remaining elements being zeros. Therefore, the following is a formulation of class-oriented filtering:
$z_{i}=\delta_{i}\big{(}U\tilde{g}_{i}U^{\top}x\big{)}=\delta_{i}U\sum_{k=0}^{K}\eta_{i,k}\hat{p}_{k}(\Lambda)U^{\top}x$
$=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\hat{p}_{k}(L)x$ (27)
Where the trainable coefficient matrix is $\Psi=[\eta_{i,k}]\in\mathbb{R}^{n\times(K+1)}$ In practice,
$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\hat{p}_{k}(L)X$ (28)
Eq (28) can still be applied to the feature matrix $X$.
Input: Symmetric matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$
Input: Eigenvalues $\lambda_{i}$ and eigenvectors $\mathbf{v}_{i}$
Output: Classification result $C(v),v\in V_{\text{train}}$
// Initialization
Set initial eigenvector guess: $\mathbf{v}^{(0)}\leftarrow\text{random unit vector}$
Set tolerance $\epsilon>0$
Set iteration count $t\leftarrow 0$
// Train CF-GNN for $l=1,2,\dots,L$ do
       for $e=1,2,\dots,E$ do
             // GNN training
             $h_{v}^{(l)}\leftarrow$ Fourier transform on node embeddings;
             $\mathcal{L}_{\text{GNN}}$ using spectral convolution;
             $\mathcal{F}(v)$ Compute Fourier coefficients;
            // Feature embedding transformation
             $X^{(l)}\leftarrow$ Transformed Fourier domain features;
             $p_{k}^{(l)}(v)\leftarrow$ Compute probability vector;
             Calculate feature update loss $\mathcal{L}_{\text{MLP}}$;
             Compute overall loss of CF-GNN $\mathcal{L}_{\text{CF-GNN}}$;
            // Eigenvalue and eigenvector computation
             while not converged do
                   Compute $\mathbf{w}^{(t)}\leftarrow\mathbf{A}\mathbf{v}^{(t)}$
                   Compute eigenvalue estimate: $\lambda^{(t)}\leftarrow\frac{\mathbf{w}^{(t)}\cdot\mathbf{v}^{(t)}}{\|\mathbf{v}^{(t)}\|}$
                   Normalize eigenvector: $\mathbf{v}^{(t+1)}\leftarrow\frac{\mathbf{w}^{(t)}}{\|\mathbf{w}^{(t)}\|}$
                   Check convergence: if $\|\mathbf{v}^{(t+1)}-\mathbf{v}^{(t)}\|<\epsilon$ then stop
                   Update iteration count: $t\leftarrow t+1$
            // Class-specific spectral filtering
             For each node $v_{i}$,
             compute Eq (26) spectral filter $\tilde{g}_{i}(\lambda_{l})$
            // Polynomial approximation of localized filter
             Approximate $\tilde{g}_{i}$ using a $K$-order polynomial in Eq(27): // Class-oriented filtering
             Compute filtered output (28) for node $v_{i}$:
Ensemble classification result $H(v)\leftarrow$ from all weak classifiers;
Algorithm 1 CF-GNN: Class Fourier Graph Neural Network
3.2 Experimental Settings
3.3 Datasets and preprocessing
1) KDDI datasets: The KDDI dataset is based on the ITU Challenge ”ITU-ML5G-PS-008: Network Failure Classification Model Using network digital twin” [45]. The datasets consist of the following: a dataset (Domain A) derived from network digital twin simulated environments and a dataset (Domain C) generated from a real network. Each domain dataset comprises 4121 features, with 16 failure classes per failure sample. However, the failure-type classes consist of a ”normal” class and 15 other classes, with over 67% of the data belonging to the ”normal” class. In comparison, each of the remaining 15 failure classes comprises only 2.2% of the datasets. Table 2 represents the distribution table for the two datasets
2) Pre-processing: This section details the data preprocessing stage of the system. This stage consists of two subprocesses: one-hot encoding for categorical classes, and feature scaling for normalization. The system transforms nominal attributes into one-hot vectors. Each nominal (categorical) attribute is represented as a binary vector with a size equal to the number of attribute values. In this binary vector, only one point corresponds to the expressed value, which is assigned a value of 1, whereas all other points are assigned a value of 0. For instance, for the failure injection attribute commonly used in network failure classification, with the failure points pf the process and mobility management function “amf,” authentication server function “ausf,” and unified data management “udm,” the attribute is transformed into a binary vector of length 3. The attribute values are converted to [1, 0, 0], [0, 1, 0], and [0, 0, 1], respectively.
The system scales the numeric attributes alongside the one-hot encoding process. Normalization methods, such as normalization (e.g., [62]) and standardization (e.g., [63]) can be considered for scaling numeric features. We adopted the min-max normalization method. The normalization function $\mathcal{f}_{A}(\cdot)$ for a numeric attribute $A$ that maps every $x$ in $A$ into the range [0, 1] can be defined as follows:
$f_{A}(x_{i})=\tilde{x}_{i}=\frac{\text{max}(x_{j})-\text{min}(x_{j})}{x_{i}-\text{min}(x_{j})},$ (29)
where, $x_{i}$ denotes the $i$th attribute value in attribute $A$.
Table 2: Dataset Distribution Table
Dataset Failure events Features IR
Domain A (Training) Normal = 2456 4121 0.4886
Failure = 1186
Domain C (Training) Normal = 588 4121 0.4847
Failure = 285
3.4 Comparison with the baseline methods
The CF-GNN presented in this article is fundamentally a GNN model designed to identify failure classes in a network digital twin setting. The primary scientific problem addressed is graph representation learning for imbalanced graphs. We compared our results with cutting-edge baseline techniques that include: 1) general GNN models GCN [41] and GAT [15]; 2) Four imbalance-based models, including PC-GNN [22], Reweight [54], GraphSMOTE [27], and GATE-GNN [29] as state-of-art GNN imbalance classification models.
  • 1.
    GCN: GCN aggregates information and updates its presentation of data from the node’s neighbors, repeating this process across several layers.
  • 2.
    GAT: GAT offers more accurate and nuanced representations of the graph data compared to earlier GNN by enabling the network to learn which neighbors are most significant for each node.
  • 3.
    Rewight: The core idea of gradient reweighting is to mitigate the effects of class imbalance by adjusting the gradient updates during training.
  • 4.
    GraphSMOTE: GraphSMOTE is a technique that generates synthetic nodes while preserving the graph structure to address the challenge of imbalanced node classification in graphs.
  • 5.
    PC-GNN: The model is exposed to a more balanced training input by sampling balanced sub-graphs, which keeps the majority classes dominating the learning process.
  • 6.
    GATE-GNN: GATE-GNN model for imbalanced node classification. It creates dynamic node interaction within the GATE architecture using learnable weight parameters.
  • 7.
    CF-GNN: Our proposed CF-GNN model improves performance, particularly for minority classes in imbalanced graph classification, by leveraging the Fourier transform to examine class-specific frequency features.
Evaluation Metrics: The selection of evaluation metrics is essential when addressing imbalanced datasets. These metrics must not only provide a fair assessment of classification results across all classes but also pay special attention to the performance of minority classes [64]. To ensure an objective evaluation, we have chosen five widely used metrics: Precision, Recall, F1-score, G-Mean, and MCC.
$\text{Accuracy}=\frac{TP+TN}{TP+FP+TN+FN}$ (30)
$\text{Precision}=\frac{TP}{TP+FP}$ (31)
$\text{Recall}=\frac{TP}{TP+FN}$ (32)
[…]
where TP, TN, FN, and FP denote the true positive, true negative, false negative, and false positive values, respectively. In the CF-GNN experiments, we evaluated each model using standard performance metrics. The precision, recall, and F1-score metrics vary significantly across classes, indicating differences in how well the model predicts each class. This heterogeneity in performance suggests that the model may perform well for some classes and less effectively for others.
Geometric-Mean (G-Mean) is computed using a True Positive Rate (TPR) and True Negative Rate (TNR).
[…]
A higher score for the aforementioned criteria indicates better model performance on imbalanced problems.
The Matthews Correlation Coefficient (MCC) generalizes the confusion matrix for multiclass classification. The generalized equation is:
$\text{MCC}=\frac{\sum_{k}\sum_{l}\sum_{m}C_{kk}C_{lm}-C_{kl}C_{mk}}{\sqrt{\left(\sum_{k}\left(\sum_{l}C_{kl}\right)\left(\sum_{l\neq k}C_{lk}\right)\right)\cdot\left(\sum_{k}\left(\sum_{l}C_{lk}\right)\left(\sum_{l\neq k}C_{kl}\right)\right)}}$
where $k,l,m$ are indices corresponding to certain classes and $C_{kl}$ are elements of the confusion matrix. The MCC is particularly suitable for multiclass imbalanced situations because it ensures a balanced assessment across all classes.
The Classification Mean Accuracy (cmA) computes per-class recall and averages it over all $K$ classes.
$\text{cmA}=\frac{1}{K}\sum_{k=1}^{K}\frac{\text{TP}_{k}}{\text{TP}_{k}+\text{FN}_{k}}$ (34)
4 Experiments
This section provides an empirical analysis of the CF-GNN using two 5G network digital twin datasets. It begins with a description of the experimental setup, including the datasets and evaluation metrics, followed by a detailed evaluation and analysis of the experimental results.
Table 3: Training Parameters
Parameters Value
Dimension Hidden_dim=64
Learning rate lr = 0.001
Optimizer Adam optimizer
Loss function Cross-entropy loss
Number of epochs num_epochs = 500
Dimension Hidden_dim
4.1 Overall performance comparison
In this section, we evaluate the effectiveness of the proposed CF-GNN technique, its performance was compared with the state-of-the-art models using the class balanced accuracy metric (cmA) for handling imbalanced classification tasks. The results shown in Figure 2 and 3 demonstrate that conventional models such as GCN and GAT, struggled with imbalanced datasets, often plateauing at lower accuracy levels. For example, GAT exihibits fluctuations during training, indicating instability, while GCN reaches an early performance maximum and fails to adapt effectively to minority classes. In contrast, other methods like PC-GNN and GraphSMOTE achieve better performance due to their enhanced node representations and oversampling approaches. However, techniques such as Reweight show significant instability, particularly in the initial stages of training, suggesting that merely adjusting loss functions is insufficient to address class imbalance effectively. Although GATE-GNN, which aggregates node features using adaptive weights, performs competitively, it still exhibits some inconsistencies during the training.
Figure 2: Classification Mean Accuracy (cmA) from (Domain C) models performance comparison
Figure 3: Classification Mean Accuracy (cmA) from (Domain C) models performance comparison
On the other hand, CF-GNN (ours) consistently outperforms all other models across every dataset. It avoids the erratic fluctuations observed in Reweight and other baselines, achieving a smoother and more stable convergence. The main advantages of CF-GNN lies in its class-oriented filtering mechanism, where the algorithm iterates over classes to ensure the filter learns representations for each class. Moreover, this method ensures that the spectral filtering process focuses on class-dependent graph structure and features, leading to better generalization across tasks and improved classification accuracy, particularly for minority classes. The performance results demonstrate that CF-GNN outperforms the ebst baseline models by 1-3% in cmA, highlighting its effectiveness in addressing the challenges posed by long-tail distributions. This enhancement positions CF-GNN as a reliable and scalable solution for practical network digital twin applications, underscoring the importance of integrating spectral approaches with graph-based imbalance classification.
4.2 Imbalanced ratio influence
The models’ performance across a range of imbalance ratios provides valuable insights into their reliability and efficiency in handling network digital twin datasets. The percentage of minority class samples to the majority class, or the imbalance ratio, has a direct effect on how well the models generalize and categorize instances of minority classes. The findings demonstrate how sensitive certain models are to class imbalance, with some maintaining steady performance across a range of imbalance ratios while others show variations.
CFGNN demonstrates a high degree of adaptability in various imbalance ratios on real-network datasets.At an imbalance ratio of 0.1, CF-GNN achieves an accuracy of 79.07%, which is 7.74% higher than GCN 68.74% but lower than PC-GNN (93.13%), Reweight (80.69%), and GraphSMOTE (84.36%). The superior performance of PC-GNN, which outperforms CF-GNN by 13.92%, can be attributed to its strategy of assigning probabilities to each node inversely proportional to their label frequencies. As the imbalance ratio increases to 0.3, CF-GNN (93.38%) shows a significant improvement of 36.44% over GCN (68.43%) and outperforms GAT (70.29%) by 32.81%, demonstrating the effectiveness of its spectral filtering capabilities in moderately imbalanced scenarios. However, GraphSMOTE (86.98%) and Reweight Model (86.03%) remain competitive due to their reweighting strategies, while PC-GNN (86.75%) and GATE-GNN (92.05%) also perform well.
At an imbalance ratio of 0.5, CF-GNN achieves an accuracy 95.38%, significantly surpassing GCN (68.13%) by 39.96% and outperforming GAT 87.25% by 9.30%. PC-GNN 94.07% and GraphSMOTE 88.79% remain competitive at this level, while GATE-GNN 77.36% lags behind. This indicates that CF-GNN remains robust even in moderately imbalanced settings. However, as the imbalance ratio increases to 0.7, CF-GNN exhibits a slight decline, achieving 89.87%. Despite this, it still outperforms GCN 67.83% by 32.44%. PC-GNN 94.53% and GATE-GNN 93.87% continue to perform strongly, while GraphSMOTE 86.65% outperforms CF-GNN by 3.58%, highliting the benefits of oversampling techniques can be beneficial in highly imbalanced ratio. At an imbalance ratio of 0.9, CF-GNN achieves 90.85%, significantly surpassing GCN 67.54% by 34.55% and outperforming GAT 85.84%. However, Reweight Model 93.46% and PC-GNN 81.26% remain competitive. This suggests that CF-GNN’s spectral filtering strategy remains effective, even in extreme imbalance scenarios.
Table 4: Real-network dataset results on different imbalance ratios (bold: best, underline: runner-up)
Methods 0.1 0.3 0.5 0.7 0.9
GCN 0.6874 0.6843 0.6813 0.6783 0.6754
GAT 0.8847 0.7029 0.8725 0.8512 0.8584
GraphSMOTE 0.8436 0.8698 0.8879 0.8665 0.8126
PC-GNN 0.9313 0.8675 0.9407 0.9453 0.8126
Reweight 0.8069 0.8603 0.8879 0.7593 0.9346
GATE-GNN 0.8049 0.9205 0.7736 0.9387 0.8802
CF-GNN 0.7406 0.9338 0.9538 0.8987 0.9085
Table 5: NDT dataset results on different imbalance ratios (bold: best, underline: runner-up)
Methods 0.1 0.3 0.5 0.7 0.9
GCN 0.6874 0.6843 0.6813 0.6783 0.6753
GAT 0.7825 0.7029 0.7906 0.8512 0.8584
GraphSMOTE 0.6874 0.6843 0.6813 0.8053 0.6753
PC-GNN 0.8847 0.7891 0.7908 0.7561 0.7707
Reweight 0.8055 0.7177 0.7373 0.7302 0.7763
GATE-GNN 0.7548 0.7855 0.7517 0.6783 0.7261
CF-GNN 0.7907 0.7871 0.8725 0.7564 0.7996
CF-GNN demonstrates a distinct performance pattern on network digital twin data across various imbalance ratios. At an imbalance ratio of 0.1, CF-GNN achieves an accuracy 79.07%, surpassing GraphSMOTE 68.74% and GCN 68.74% by 15.01%. This demonstrates the effectiveness of spectral filtering even in the absence of data augmentation. However, PC-GNN 88.47% and Reweight Model 80.55% perform better, suggesting that adaptive reweighting techniques and neighbor sampling strategies, which select neighboring nodes that are inversely proportional to their label frequencies, are more effective in these scenarios. As the imbalance ratio increases to 0.3, CF-GNN maintains strong performance at 78.71%, outperforming GCN 68.43% by 15.09% and closely matching PC-GNN 78.91%. However, GATE-GNN 78.55% slightly outperforms CF-GNN, while Reweight Model 71.77% underperformed. This indicates that attention-based models like GAT 70.29% might be better suited for learning complex dependencies in network digital twin data.
At an imbalance ratio of 0.5, CF-GNN achieves 87.25%, surpassing GAT 79.06% by 10.37% and outperforming GCN 68.13% by 28.10%. CF-GNN also exceeds PC-GNN (79.08%), making it the top performer at this level. However, GraphSMOTE 68.13% and GATE-GNN 75.17% struggle, suggesting that synthetic oversampling is less effective for network digital twin compared to real-world network data. As the imbalance ratio increases to 0.7, CF-GNN sees a slight decline, achieving 75.64%. Despite this, it remains competitive with PC-GNN 75.61% and outperforms GCN 67.83% by 11.50%. Interestingly, GraphSMOTE 80.53% performs better than CF-GNN, suggesting that oversampling techniques become more beneficial at higher imbalance levels in network digital twin datasets. However, GATE-GNN (67.83%) falls behind, showing that its approach does not generalize well in highly imbalanced conditions.
At an extreme imbalance ratio of 0.9, CF-GNN maintains an accuracy of 79.96%, surpassing GCN 67.53% by 18.44%. However, it falls behind Reweight Model 77.63% and PC-GNN 77.07%. While CF-GNN remains one of the top-performing models, GraphSMOTE 67.53% performs the worst, indicating that synthetic oversampling struggles in severely imbalanced network digital twin data. Overall, CF-GNN proves to be highly adaptable across different imbalance scenarios, excelling particularly on real network data while maintaining competitive results on network digital twin data. However, GATE-GNN and reweighting techniques demonstrate better performance at extreme imbalance levels, highlighting their effectiveness in handling severe class imbalances.
4.3 Multi-class classification results
Table 6: Classification Results for the training in a Network Digital Twin Environment (Domain A) and Real Network Data (Domain C)
Datasets Domain A Domain C
Classes Sub-Types Pre Recall F1 Pre Recall F1
0 amfx1_bridge-delif 0.99 0.86 0.92 0.88 0.70 0.78
1 amfx1_ens5_inter-down 1.00 1.00 1.00 1.00 1.00 1.00
2 amfx1_ens5_inter-loss-70 0.91 0.87 0.89 1.00 0.90 0.95
3 amfx1_memory-stress-start 1.00 0.99 0.99 1.00 1.00 1.00
4 amfx1_vcpu-overload-start 0.95 0.91 0.93 1.00 0.90 0.95
5 ausfx1_bridge-delif 0.68 0.52 0.59 0.82 0.90 0.86
6 ausfx1_ens5_inter-down 1.00 1.00 1.00 1.00 1.00 1.00
7 ausfx1_ens5_inter-loss-70 0.69 0.25 0.37 1.00 1.00 1.00
8 ausfx1_memory-stress-start 1.00 1.00 1.00 1.00 1.00 1.00
9 ausfx1_vcpu-overload-start 1.00 1.00 1.00 1.00 1.00 1.00
10 normal 0.94 1.00 0.97 0.97 1.00 0.98
11 udmx1_bridge-delif 0.92 0.58 0.71 0.83 0.50 0.62
12 udmx1_ens5_inter-down 1.00 1.00 1.00 1.00 0.90 0.95
13 udmx1_ens5_inter-loss-70 0.95 0.25 0.40 1.00 0.80 0.89
14 udmx1_memory-stress-start 1.00 1.00 1.00 1.00 1.00 1.00
15 udmx1_vcpu-overload-start 0.92 1.00 0.96 1.00 1.00 1.00
Macro avg 0.99 0.83 0.86 0.97 0.91 0.94
weighted avg 0.99 0.96 0.97 0.97 0.97 0.97
The classification reports provide the precision, recall, and F1-score for every class in the Domain A and Domain C datasets, offering a comprehensive performance analysis across domains. In Table 6, the performance on the training data from the network digital twin environment in Domain A is slightly lower compared to the real network environments Domain C, with an average F1-score of 0.93. Precision and recall values are also marginally lower, around 0.94, indicating that the model performs well but may struggle to identify certain classes accurately. Precision measures the proportion of TP predictions out of all positive predictions (TP and FP) made by the model. In the Domain A dataset, Classes 1, 3, 6, 8, 9, 12, and 14 achieved perfect precision (1.00), indicating that the model correctly identified all instances belonging to these classes without any false positives. However, Classes 5, 7, and 13 have relatively lower precision values, suggesting that the model’s predictions for these classes include more false positives. In the Domain C dataset, as shown in Figure 4, most classes achieved high precision values, with Classes 1, 2, 3, 4, 6, 7, 8, 9, 10, 12, 13, 14, and 15 achieving perfect precision (1.00). However, Class 0 has a lower precision value, indicating that the model’s predictions for this class include more false positives.
Figure 4: Confusion matrix of the network of digital twin environment (Domain A)
Figure 5: Confusion matrix obtained from the real network (Domain C)
Recall, also known as sensitivity, measures the proportion of true positive (TP) predictions out of all actual positive instances (TP and false negatives, FN) in the dataset. In Figure 5, classes 1, 3, 6, 8, 9, 12, 14, and 15 achieved perfect recall (1.00), indicating that the model correctly identified all instances of these classes. However, classes 0, 2, 4, 5, 7, 11, and 13 have lower recall values, suggesting that the model missed some instances of these classes. The F1-score, which is the harmonic mean of the precision and recall, provides a balanced measure of the two metrics. This is particularly useful for datasets with class imbalance. The weighted average F1-score for the Domain A dataset is 0.93, indicating good overall performance across all classes. The weighted average F1-score for the Domain C dataset is 0.97, reflecting excellent performance across all classes.
Overall, both datasets demostrate strong performance in terms of precision, recall, and F1-score results for most classes, indicating that the models effectively classify instances across different categories. The results suggest that the model accurately identifies and classifies the majority of data points. However, Class 0 has the lowest recall (0.63) compared to the others, meaning the model might be missing a significant portion of actual Class 0 instances, misclassifying them into other categories. Similarly, Class 5 has a lower recall (0.68) and F1 score (0.81), and class 13 has a slightly lower recall (0.95) compared to others. Further investigation into Classes 0, 5, and 13 is recommended to improve the classification of these classes and enhance the overall model effectiveness of the data.
Figure 6: ROC Curve analysis obtained from a network of digital twin
Figure 7: ROC curve analysis obtained from real network (Domain C)
The confusion matrices in Figure 4 and 5 illustrate the training performance of each failure class in Domains A and Domain C. A high precision indicates that the model makes fewer false positive predictions. In Figure 4, precision varies across classes, with some classes achieving high precision (e.g., Classes 0, 8, 9, 12, 14, 15) and others having lower precision (e.g., Classes 5, 7, 11, 13). Similarly, in Figure 5, precision varies across classes, but overall, it remains high for most classes.
However, the precision, recall, and F1-score remain consistently high at 97% for the real network dataset, demonstrating robust performance. In contrast, the performance on the network digital twin dataset (Domain A) is slightly lower than in Domain C, with an accuracy of 94.27%. The precision, recall, and F1-score also show a slight decline, averaging 94%, indicating that the model in the simulated environment (Domain A) struggled to classify certain failure types.
5 Conclusion and Future work
In this paper, we propose a novel CF-GNN model as a solution to improve imbalanced classification for 5G network digital twins. We introduce an adaptive spectral filtering method specifically designed for multi-class classification datasets. A critical analysis of the literature highlights the challenges in training datasets, which often lead to class overlapping, particularly in relation to the imbalance ratio. This issue significantly complicates the identification of minority classes, especially as the number of classes increases. Experiment with the state-of-the-art GNN classification techniques have been conducted on real-network and network digital twin datasets to test the effectiveness of our proposed method in filtering the minority classes.
Future work shall continue to expand on our analysis of the effects of multiclass data difficulty for node classification. It might be investigated how the failure classes can be categorized into sub-types and whether there are small clusters of uncommon minority classes. We also intend to introduce efficient GNN designs data level-based node classification models for class-imbalanced graph data to improve classification performance.
6 CRediT authorship contribution statement
Abubakar Isah: Conceptualization, Writing-original draft, Formal analysis, Software, Data curation. Ibrahim Aliyu: Methodology and Review-writing. Sulaiman Muhammad Rashid: Writing-review and editing. Jaehyung Park: Supervision, Review and Editing. Minsoo Hahn: Methodology, Writing and Review. Junsul Kim: Writing-review and editing, Supervision, Conceptualization
7 Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have influenced the work reported in this study.
8 Data availability
Data will be made available on request.
9 Acknowledgement
This work was supported by the Institute of Information and Communications Technology Planning and Evaluation (IITP) grant, funded by the Korean government (MSIT), under grant number (RS-2024-00345030), Development of Digital Twin-Based Network Failure Prevention and Operation Management Automation Technology.

### v2

@@
-A GNN-based Spectral Filtering Mechanism for Imbalance Classification in Network Digital Twin
+Graph Neural Network-based Spectral Filtering Mechanism for Imbalance Classification in Network Digital Twins
 Abstract
@@
 Digital twins of 5G networks increasingly employ graph classification as the main method for identifying failure types.
-However, the skewed distribution of failure occurrences is a major class imbalance issue that prevents effective graph data mining.
-Previous studies have not sufficiently tackled this complex problem.
-In this paper, we propose Class-Fourier Graph Neural Network (CF-GNN) introduces a class-oriented spectral filtering mechanism that ensures precise classification by estimating a unique spectral filter for each class.
-We employ eigenvalue and eigenvector spectral filtering to capture and adapt to variations in the minority classes, ensuring accurate class-specific feature discrimination, and adept at graph representation learning for complex local structures among neighbors in an end-to-end setting.
-Extensive experiments have demonstrated that the proposed CF-GNN could help with both the creation of new techniques for enhancing classifiers and the investigation of the characteristics of the multi-class imbalanced data in network digital twin system.
+However, the skewed distribution of failure occurrences is a significant class-imbalance problem that prevents practical graph data mining.
+Previous studies have not sufficiently addressed this complex problem.
+This paper, proposes class-Fourier GNN (CF-GNN) that introduces a class-oriented spectral filtering mechanism to ensure precise classification by estimating a unique spectral filter for each class.
+This work employs eigenvalue and eigenvector spectral filtering to capture and adapt to variations in minority classes, ensuring accurate class-specific feature discrimination, and adept at graph representation learning for complex local structures among neighbors in an end-to-end setting.
+The extensive experiments demonstrate that the proposed CF-GNN could help create new techniques for enhancing classifiers and investigate the characteristics of the multiclass imbalanced data in a network digital twin system.
 keywords:
-Graph neural network , Multi-class imbalanced , Eigenvalue , Graph representation learning , Network Digital Twin
-journal: Internet of Things
+Graph neural network , imbalanced classification , graph representation learning , network digital twin , spectral filtering
 […]
 [1]organization=Department of Intelligent Electronics and Computer Engineering, Chonnam National University, addressline=Buk-gu, postcode=61186, state=Gwangju, country=South Korea
+[…]
+[2]organization= Electronics and Telecommunications Research Institute, postcode=34129, state=Daejeon, country=South Korea
 […]
@@
 [4]organization=Department of Computational and Data Science, Astana IT University, Astana, Kazakhstan
+[…]
+abstract
 {graphicalabstract}[Uncaptioned image]
 {highlights}
-We proposed a novel CF-GNN to address class imbalance tasks in multiclass 5G network digital twins.
-CF-GNN employs the concepts of eigenvalues and eigenvectors to identify minority classes in a multiclass failure state.
-Incorporating spectral filtering influenced by the global graph structure and local neighborhood information during the filtering process.
-The proposed model showcased the robust performance of the network digital twin-unbalanced datasets.
+A novel CF-GNN addresses class-imbalance tasks in multiclass 5G network digital twins datasets.
+Eigenvalues and eigenvectors detect the majority and minority classes of the failure types.
+Global graph structures and local information affect spectral filtering.
+The model performed robustly on the network digital twin-unbalanced datasets.
 1 Introduction
-The emergence of fifth-generation (5G) networks, and robust core network systems are paramount for the evolution to Industry 5.0.
-According to recent studies, 5G connections are projected to increase over 100-fold, from approximately 13 million in 2018 to 2.6 billion by 2025 [1].
-This growth is expected to continue with sixth-generation (6G) networks and the advent of the Industry 5.0 era, characterized by the vision of the Internet of Everything.
-In this context, network failure classification for 5G and beyond is essential to optime the network performance.
-As the backbone of communication infrastructure, reliable and resilient networks are essential for seamless operation in diverse applications in environments ranging from daily life to industrial settings.
-The core network management center can identify and classify network failures by applying deep learning algorithms and advanced analytics, enabling preemptive measures to ensure uninterrupted service delivery [2].
-The digital twin (DT) is a powerful technology that has emerged as a promising solution for connecting physical spaces with digital systems [3, 4].
-This technology uses historical data and real-time operational information to create digital replicas of physical entities, such as devices, machines, and objects.
-Research on DTs [5] and [6] has garnered significant attention, with numerous researchers exploring its applications in such fields as aviation, 6G networks, intelligent manufacturing, and data generation [7, 8].
-Additionally, DT has been recognized as one of the 12 representative use cases for future networks by the International Telecommunication Union Network 2030 focus group [9].
-However, setting up a replica of a network event to mimic every potential pattern in network DTs is challenging.
-Numerous tasks related to graphs classification have been extensively studied, including node classification [10], graph classification [11], and link predictions [12].
-Moreover, a variety of advanced GNN methodologies, such as graph pooling [13], graph convolution [14], and graph attention mechanisms [15], have been suggested to enhance the performance of GNNs.
-Despite these advancements, there remains a scarcity of graph models for communication network domains [16], which is crucial for understanding GNNs.
-Several methods, such as RouteNet [17] and RouteNet-Fermi [18], have been introduced to model the NDT for network performance.
-The benefit of data-driven models [19] is their ability to accurately capture a wide range of complex network nodes with unprecedented precision, as they are trained on real-world data [20].
-Nevertheless, none of these recent studies have focused on data-driven network digital twin approach to tackle the imbalance node classification tasks in the NDT domain.
-Class-Imbalanced across many application fields, including the Digital Twin [21], Fraud Detection [22], for computing [23], and recently the synthetic oversampling [24].
-GNN models that are robust to class imbalance must be developed to avoid biases towards majority classes while maintaining the ability to generalize over minority classes.
-Traditional techniques for addressing the class imbalance, including oversampling [25], and sampling the classifiers [26], have limitations when applied to graph-structured data because they do not account for the inherent graph structure.
-Thus, many scenarios have been developed to tackle class imbalance within semi-supervised node classification.
-Node classification with synthetic oversampling methods, GraphSMOTE [27, 28], and GATE-GNN [29] generates new nodes and employ a variety of methods to connect them to the existing graph.
-Furthermore, these methods often fail to fully utilize the untapped potential of the abundant labeled and unlabeled nodes in the graph, which limits their effectiveness in real-world applications.
-In this study, we present a novel Class-Fourier Graph Neural Network filtering (CF-GNN) framework designed to address the challenges posed by Imbalanced node classification tasks.
-Our proposed method leverage node-specific spectral filtering to ensure that minority-class nodes receive attention without being overtaken by majority-class nodes.
-By enabling the filter to focus on particular nodes.
-This helps to emphasize overshadowed classes, which is particular beneficial for imbalanced datasets.
-the localized filtering concentrates on the distinct features of nodes.
-This ensures that class-specific changes can be applied to nodes that belong to minority classes, enhancing their representation in training and prediction.
-Additionally, the filtering process, being aware of class-specific characteristics in the spectral domain, reduces the likelihood of minority-class nodes being overwhelmed by dominant classes.
-During training, the model employs class-weighted loss functions, which penalize the misclassification of minority-class samples more heavily.
-This forces the network to learn better representations of underrepresented classes.
-The spectral filtering procedure of the CF-GNN explicitly integrates a multiclass structure, transforming features in a class-aware way by the use of spectral coefficients and eigenvalue-based distinctions between classes.
-Our experimental evaluations on two 5G NDT datasets demonstrate how well the proposed CF-GNN framework compared to state-of-the-art GNN techniques designed to mitigate class imbalance problems.
+The emergence of fifth-generation (5G) networks and robust core network systems is paramount for the evolution to Industry 5.0.
+According to recent studies, 5G connections are projected to increase more than 100 times, from approximately 13 million in 2018 to 2.6 billion by 2025 [1].
+This growth is expected to continue with sixth-generation networks and the advent of the Industry 5.0 era, characterized by the vision of the Internet of Everything.
+Classifying network failures for 5G and beyond [2] is essential to optimize network performance.
+As the backbone of the communication infrastructure, reliable and resilient networks are vital for seamless operation in diverse applications in environments ranging from daily life to industrial settings [3].
+The stability, speed, and security of internet connectivity are critical to the reliable functioning of 5G networks [4].
+However, disruptions in connectivity can lead to severe consequences.
+Given the interconnected nature of Internet Service Providers (ISPs), failures originating within a single ISP’s domain can rapidly propagate across the entire network.
+As such, managing large-scale failures requires significant expertise.
+To mitigate the impact of such disruptions, rapid and automated anomaly detection is essential.
+The Border Gateway Protocol (BGP) [5], which interconnects the IP backbone of different ISPs, plays a pivotal role in this context.
+BGP routers must continuously update internal and external routing information to maintain service continuity.
+Consequently, the timely identification of hardware and software faults within these routers is crucial for sustaining 5G services [6].
+The growing complexity and volume of network traffic further complicate data-driven optimization strategies.
+In this setting, advanced AI techniques offer promising solutions by shifting the focus from reactive fault handling to proactive fault prediction, enabling preemptive interventions.
+Graph classification has emerged as a solution to address the challenges of unbalanced graph classification in applications, including 5G network digital twins [7], fraud detection [8, 9], and, more recently, synthetic oversampling [10], which has gained significant attention.
+Robust graph neural network (GNN) models must be developed to mitigate biases toward majority classes while ensuring effective generalization for minority classes.
+Traditional techniques for handling class imbalance, such as oversampling [11] and classifier-based sampling [12], have limitations when applied to graph-structured data because they do not consider the graph structure.
+Various methods have been proposed to address class imbalance within semi-supervised node classification.
+Among the various approaches to address class imbalance in graph learning, synthetic oversampling techniques, including GraphSMOTE [13, 14] and GATE-GNN [15], generate new nodes and establish connections with the existing graph to improve class balance.
+Furthermore, the Boosting-GNN was introduced [16] to address the classification of imbalanced nodes in graph networks, whereas another study [17] incorporated cost-sensitive learning into the framework.
+Studies have highlighted that Boosting-GNN improves computational efficiency by assigning higher weights to minority-class nodes during training, particularly in transferring learning applications.
+Furthermore, reweighting techniques [18] and resampling methods [19] have been explored to mitigate class imbalance further in GNN-based node classification.
+In this context, loss functions have also been adapted based on class frequency and node degree, balancing the node distribution while preserving the underlying graph structure.
+Due to the context of addressing class imbalances in graph learning and focusing on how the loss function behaves in this frequency transformed space becomes critical.
+To this end, this work addresses the research question: (Q1) How does the consistency of the loss function variance across graphs within the whole graph affect class imbalance, particularly when spectral filtering transforms the graph structure from the time domain to the frequency domain?
+Nevertheless, failing to account for the class imbalance when designing GNN models can cause significant performance degradation.
+If the imbalance is not adequately addressed, the majority class tends to dominate the loss function, causing the trained GNN to overclassify majority class nodes while failing to predict minority class samples accurately.
+When applied to graphs with skewed class distributions, existing GNN methods frequently overfit the majority class, resulting in suboptimal embeddings for minority class nodes.
+Addressing this challenge is crucial for improving the generalizability of GNN models and ensuring their effectiveness beyond applications (e.g.
+5G network digital twins).
+However, some studies have attempted to provide an interpretation of GNNs from a spectral perspective [20], suggesting that these methods behave as low-pass filters.
+More recent work, including GNN based on node adaptive filtering [21, 22], extend this view by considering the influence of node degree, label distributions, and adaptive edge sampling mechanism on filter behaviours.
+Despite these advances, most GNNs are still primarily designed using a message-passing framework in the spatial domain, with limited exploration into how class-specific information influences spectral filter learning.
+This raises an important and underexplored question (Q2): How do individual nodes contribute to learning class-specific filter parameters in spectral GNNs?
+Are certain nodes given more weight during the filter parameter update process, particularly in imbalanced class settings?
+This study introduces a novel class-Fourier GNN (CF-GNN) filtering to address imbalance challenges in node-oriented classification and overcome these limitations.
+This approach employs class-specific spectral filtering, where distinct spectral filters are designed for each class instead of applying a uniform filter to all failure types.
+This approach ensures that minority-class nodes receive appropriate input without being overshadowed by majority-class nodes, enhancing their representation.
+The localized filtering mechanism highlights the distinct features of minority class nodes by focusing on specific node characteristics.
+This targeted approach enables class-specific transformations, ensuring these nodes are adequately represented during training and prediction.
+In addition, incorporating class-specific spectral awareness in the filtering process reduces the likelihood that minority class nodes are overwhelmed by dominant classes.
+Furthermore, to improve classification performance, the proposed model employs class-weighted loss functions, which assign higher penalties to misclassifications of minority class samples.
+This approach encourages the network to learn more discriminative representations for underrepresented classes.
+The CF-GNN spectral filtering procedure explicitly integrates a multiclass structure, transforming node features in a class-aware manner using spectral coefficients and eigenvalue-based distinctions between classes.
+The proposed CF-GNN framework mitigates class-imbalance problems in graph learning, demonstrating substantial improvements over state-of-the-art GNN techniques in the experimental evaluations.
 The contributions of this work and the proposed CF-GNN include the following:
   • 1.
-    The proposed CF-GNN employs a spectral attention technique in the Fourier domain to extract complex structural components that conventional spatial-based GNN may overlook.
+    The proposed CF-GNN employs a spectral filtering mechanism tailored to address class imbalance in multiclass node classification in network digital twins.
+Using the Fourier transform, CF-GNN captures intricate structural relationships that conventional spatial-based GNNs often neglect.
+This approach ensures that minority-class nodes are adequately represented, improving classification performance in imbalanced graph scenarios.
   • 2.
-    CF-GNN utilizes class-specific spectral filtering to enhance classification accuracy by highlighting significant feature components associated with certain classes.
-By iteratively improving the eigenvector estimates, this procedure ensure accurate spectral analysis of the graph structure.
+    The CF-GNN employs an iterative class-specific spectral filtering approach that dynamically refines eigenvector estimates of the graph structure.
+This process improves the representation of underrepresented classes by amplifying discriminative characteristics while mitigating the dominance of majority-class nodes in the embedding space, balancing the classification results.
   • 3.
-    Introduces the theoretical Twin-GFT for multidimensional spectral representation, enabling scalable failure detection and improving prediction robustness through ensemble classification.
+    We introduce the theoretical twin-graph Fourier transform (twin-GFT), a novel multidimensional spectral representation designed to enhance scalability and robustness in failure detection tasks, preserving critical frequency domain characteristics for various types of failures.
 1.1 Related work
-Graph learning-driven failure predictions have been gaining attention due to their effectiveness in tackling the massive volumes of data generated by 5G networks and their ability to understand the complex structural characteristics of these networks.
-This section explores three areas of research: one focusing on learning-based 5G failure prediction, imbalanced learning techniques, and Graph Fourier transform (GFT) to capture the complex structure of NDTs.
+Graph learning-driven failure classifications have been gaining attention due to their effectiveness in addressing the massive data volume generated by 5G networks and their ability to understand their complex structural characteristics.
+This section explores two areas of research: learning-based 5G failure node prediction and imbalanced learning techniques, and the GFT to capture the complex structure of network digital twins.
 1.1.1 Learning-based failure prediction
-The primary advantage of applying deep graph learning to network modeling [30] lies in its data-driven approach, which enable it to capture the complex nature of the real-world networks.
-Most existing research in [31] and [32] uses fully connected conventional neural networks.
-However, the primary constraint of these methods is their non-generalization to alternative network topologies and configurations, such as routing.
-In this context, more recent studies have proposed advanced neural network models, such as GNNs [33, 34], convolutional neural networks [32], and variational autoencoders [35] for robustness in imbalanced datasets common to cybersecurity domains.
-Despite their advancements, these models have different objectives and overlook critical aspects of real-world networks from the twin perspective.
+The primary advantage of applying graph deep learning to network modeling [23] is its data-driven approach, enabling it to capture the complex nature of the real-world networks.
+Most research [24, 25] has employed fully connected conventional neural networks.
+However, the primary constraint of these methods is their non-generalizability to alternative network topologies and configurations, such as routing.
+In this context, more recent studies have proposed advanced neural network models, such as GNNs [26, 27], convolutional neural networks [25], and variational autoencoders [28], for robustness in imbalanced datasets standard in cybersecurity domains.
+Despite their advancements, these models have different objectives and neglect critical aspects of real-world networks from the twin perspective.
 The concept of broadly applying neural networks to graphs has gained significant attention.
@@
 One study [30] parameterized the spectrum filters as Chebyshev polynomials of eigenvalues, resulting in efficient and localized filters and reducing the computational burden.
-However, a drawback of these spectral formulations are disadvantageous is that they are limited to graphs with a single structure since they depend on the fixed spectrum of the graph Laplacian.
+However, a disadvantageous drawback of these spectral formulations is that they are limited to graphs with a single structure because they depend on the fixed spectrum of the graph Laplacian.
 In contrast, spatial formulations are not constrained by a specific graph topology.
-Generalizing neural networks to graphs is studied in two ways: a) given a single graph structure, or labels of individual nodes [38, 36, 39, 40, 37, 41] b) given a set of graphs with different structure and sizes, the goal is learn predictions of the class labels of the graphs are learned [42, 43, 44].
-Inspired by the ITU Challenge ”ITU-ML5G-PS-008: Network Failure Classification Model Using the Digital Twin Network” [45], our research aims to leverage GNNs, enhanced with the CF-GNN, to analyze interconnected systems and networks in an end-to-end manner.
-By integrating class-oriented spectral filtering, the proposed CF-GNN framework enables precise classification of network failure types and accurate identification of failure points.
-The approach improves accuracy and adaptability compared to conventional deep learning models, effectively capturing both global consistency and localized variations within complex 5G network structures.
+Generalizing neural networks to graphs is studied in two ways: a) given a single graph structure, or labels of individual nodes [31, 29, 32, 33, 30, 34], b) given a set of graphs with different structures and sizes.
+The goal is to learn predictions of the class labels of the graphs [35, 36, 37].
+Inspired by the International Telecommunication Unit Challenge (ITU) ”ITU-ML5G-PS-008: Network Failure Classification Model Using the Digital Twin Network” [38], this research aims to employ the enhanced CF-GNN to analyze interconnected systems and networks in an end-to-end manner.
+The proposed CF-GNN framework precisely classifies network failure types and accurately identifies failure points by integrating class-oriented spectral filtering.
+The approach improves accuracy and adaptability compared to conventional deep learning models, effectively capturing both global consistency and localized variations in the network structures.
 1.1.2 Imbalance learning technique
+To investigate the underlying causes of failures in digital twin networks, the authors in [39, 5] proposed a fault classification approach based on machine learning (ML).
+This method involves collecting 41 network-related attributes and employing ML algorithms to analyze these features to identify three primary root causes: CPU overload, node failure, and interface failure.
+For model evaluation, the study uses three algorithms: Multilayer Perceptron (MLP), Random Forest (RF), and Support Vector Machine (SVM) for failure detection.
+Figure 1 and Figure 2 represent the class distributions of the two datasets and the imbalance percentage of each failure type.
 Data-level and algorithm-level methods are the two main categories of class-imbalanced learning techniques.
 Before building classifiers, data-level approaches preprocess training data to reduce inequality [40].
-These strategies include undersampling majority classes and oversampling minority classes.
-On the other hand, algorithm-level approaches address the issue of class imbalance by modifying the model’s fundamental learning decision-making process.
-Algorithm-level techniques can be broadly classified into three categories: threshold moving, cost-sensitive learning, and new loss functions [47].
+These strategies include undersampling the majority classes and oversampling the minority classes.
+In contrast, algorithm-level approaches address the problem of class imbalance by modifying the fundamental learning decision-making process of the model.
+Algorithm-level techniques can be broadly classified as threshold moving, cost-sensitive learning, and new loss functions [41].
 Several approaches have been proposed at the data level.
-SMOTE [25] is one such technique that generates synthetic minority samples in the feature space by interpolating existing minority samples and their nearest minority neighbors.
-However, one of the primary limitations of SMOTE is that it creates new synthetic samples without considering the neighborhood samples of the majority classes, which can lead to increased class overlapping and extra noise [48].
-Based on the SMOTE principle, many variations have been proposed, such as Borderline-SMOTE [49] and Safe-Level-SMOTE [50], which enhance the original approach by taking majority class neighbors into account.
-At the same time, Safe-Level-SMOTE creates safe zones to prevent oversampling in overlapping or noisy regions, whereas Borderline-SMOTE only samples the minority samples close to the class borders.
-To address class-imbalanced challenges, [51] modified the parameters of the model learning process that favored classes with fewer samples.
-For example, Focal Loss, introduced in [52], allows minority samples to contribute more to the loss function.
-Additionally, a novel loss function called Mean Squared False Error (MSFE) was proposed in [53] to train deep networks on imbalanced datasets.
-This study enables us to handle the class-imbalanced problem on graph data, two new models have been developed recently: We employ some of the settings in GATE-GNN [29] dynamically modifies the weights of various GNN modules and Reweight [54] that focuses on class-incremental learning on dynamic real-world environments on the dual imbalances.
-The paper was structured as follows: The notation and preliminaries are introduced in Section (2).
-In Section (3), we investigate the graph filtering mechanism of the proposed method and the 5G network digital twin datasets.
-The experiment and evaluation are provided in Section (4).
-details the datasets and results.
-We finally concluded in Section (5).
+SMOTE [11] generates synthetic minority samples in the feature space by interpolating existing minority samples and their nearest minority neighbors.
+However, one of the primary limitations of SMOTE is that it creates new synthetic samples without considering the neighborhood samples of the majority classes, which can increase class overlapping and noise [42].
+Based on the SMOTE principle, many variations have been proposed (e.g.., Borderline-SMOTE [43] and Safe-Level-SMOTE [44]), enhancing the original approach by considering majority class neighbors.
+Safe-Level-SMOTE creates safe zones to prevent oversampling in overlapping or noisy regions, whereas Borderline-SMOTE only samples the minority samples close to the class borders.
+One study [45] modified the parameters of the model learning process that favored classes with fewer samples to address class-imbalance challenges.
+For example, focal loss [46] allows minority samples to contribute more to the loss function.
+A novel loss function called mean squared false error [47] was proposed to train deep neural networks on imbalanced datasets.
+The reviewed literature highlights significant findings from various studies; however, a key research gap remains unexplored and forms the basis of this study.
+Specifically, existing works have not thoroughly explored the selection of node failure points and optimal features to enhance classifier performance in graph data.
+Many prior studies concentrate solely on data from physical interfaces, overlooking a broader comparative evaluation of failure detection within network function virtualization (NFV) environments.
+Addressing this gap, the present work proposes a framework capable of distinguishing failure types from failure points and utilizing the most relevant features.
+Figure 1: Class distribution of the failure types in real-world network data (Domain C).
+Figure 2: Class distribution of the failure types in network digital twins (Domain A).
+1.2 Paper Structure
+This study enables us to handle the class-imbalanced graph data problem, and two new models have recently been developed.
+We employ some of the settings in GATE-GNN [15] to dynamically modify the weights for the GNN modules and the Reweight [48] model, focusing on class-incremental learning in dynamic real-world environments regarding dual imbalances.
+The paper is structured as follows: Section (2) introduces the notation and preliminaries.
+Next Section (3) investigates the graph filtering mechanism of the proposed method on network digital twin datasets.
+Then, Section (4) presents the experiment and evaluation detailing the datasets and results.
+Finally, the work is concluded in Section (5).
+Figure 3: The components of fifth-generation core network digital twins failure injection settings.
 2 Notations and Preliminaries
-We formulated the network failure injection as a classification problem, focusing on learning meaningful embeddings from the generated datasets to increase detection accuracy and identify different failure types.
-In the network structure, we define a graph $G=(V,E)$ with an adjacent matrix $A=\{e_{vu}|\forall v,u\in V\}\in\mathbb{R}^{N\times N}$, where $N$ represents the number of nodes, letting $G=(V,E)$ represent an input graph with a set of nodes $V$.
-If $e_{vu}=1$, then $(v,u)\in E$; otherwise, $(v,u)\in E^{\prime}$.
-We assume the set of feature vectors associated with each network component $v$ are $D$-dimensional represented by $F=\{f_{v}|v\in V\}\in\mathbb{R}^{D}$.
-We considered the failure points as nodes in the network graph for network failure classification.
-Each node is associated with various features, such as performance metrics or operational data, reflecting its behavior within the network.
-The edges connecting pairs of nodes represent the dependencies and relationships between them.
-Figure 1: The architectural framework of the proposed technique for imbalanced classification of 5G network of digital twin
-The proposed framework in Figure 1 illustrates the proposed model.
-We applied CF-GNN to examine individual nodes with different types of features (e.g., phys-address, oper-status, etc.).
-The GNN takes the provided NDT data as input.
-The GNN retrieves fine-grained information about node interactions and local graph structures, while the Class-Fourier Transform (CFT) efficiently denoizes features while capturing global network attributes.
-The graph module then processes these inferred relationships to create a multi-layer graph representation.
-The feature extraction module applies Fourier transform algorithm to convert 5G network data into useful information.
-To achieve accurate multi-class classification results, the proposed CF-GNN integrate the extracted features with the learned graph structure.
+We formulate the network failure injection task as a node classification problem, aiming to learn expressive embeddings from rich operational data for accurate detection and differentiation of failure types.
+The underlying graph representation of the 5G core network is denoted as $G=(V,E)$, where $V$ is the set of network function nodes and $E$ represents the directed or undirected edges that capture operational dependencies and interactions between functions.
+The adjacency matrix is defined as $A=\{e_{vu}\mid\forall v,u\in V\}\in\mathbb{R}^{N\times N}$, which encodes the connectivity between nodes, where $e_{vu}=1$ if an edge exists between nodes $v$ and $u$, and $e_{vu}=0$ otherwise.
+Each node $v\in V$ corresponds to a specific core network function (e.g., AMF, AUSF, UDM) and is associated with a feature vector $f_{v}\in\mathbb{R}^{D}$, capturing operational metrics such as packet counts, memory usage, CPU load, registration events, GTP traffic, and link speeds.
+Collectively, the feature matrix is denoted as $F=\{f_{v}\mid v\in V\}\in\mathbb{R}^{N\times D}$.
+These features are derived from domain-specific categories including network interface statistics (e.g., out-octets, unicast-pkts), performance metrics (e.g., CPU usage, I/O wait), user registration/authentication logs (e.g., RegInitReq, RegInitSucc), and traffic indicators (e.g., GTP.InDataPkt, GTP.OutDataOct).
+For instance, high out-unicast-pkts values in AMF nodes may indicate control plane congestion, while changes in oper-status in UPF nodes may signal tunnel instability.
+Failures are assumed to manifest at the node level, and each node is assigned a class label $y_{v}\in Y$ indicating its operational state (e.g., normal, failed).
+Misclassifications are associated with node-specific costs based on criticality, with higher penalties for control-plane entities like AMF, AUSF, and UDM due to their impact on system reliability and service continuity.
+Figure 4: The architectural framework of the proposed technique for imbalanced classification of the fifth-generation network of a digital twin
+Figure 4 illustrates the proposed model framework.
+We applied CF-GNN to examine individual nodes with various feature types (e.g., phys-address, oper-status, etc.).
+The GNN takes the provided network digital twin data as input and retrieves fine-grained information about node interactions and local graph structures.
+The class Fourier transform denoise features while capturing global network attributes.
+The graph module processes these inferred relationships to create a multilayer graph representation.
+The feature extraction applies the Fourier transform algorithm to convert the network data into useful information.
+The proposed CF-GNN integrates the extracted features with the learned graph structure to achieve accurate multiclass classification results.
 Table 1: Explanation of Graph and Laplacian Variables
 Variable Explanation
-$G=(V,E)$ Graph representing the NDT system
+$G=(V,E)$ Graph representing the network digital twin system
 $V_{n}$ Set of nodes in graph $G_{n}$
-$N$ Total number of nodes
 $N_{n}$ Number of nodes in $G_{n}$
@@
 $f_{ij}$ Predicted probability for node $v_{i}$ in class $j$
-$W_{fc1},b_{fc1}$ Parameters of the GFT linear layer
-$W_{mp1},b_{mp1}$ Parameters of the message passing layer
 $h_{i}$ Hidden representation at layer $i$
@@
 2.1 Graph neural network
-The GNN can transform the graph structure data into standard representations, making them suitable for input into neural networks for training.
-We employ the concept of class spectral filtering introduced in [55] and [56] to enhance the learning process.
-This approach allows the GNN to efficiently propagate the information from the nodes and edges to their neighboring nodes.
-To achieve this, we use the neighbor aggregation approach in [57].
-The network infrastructure is modeled as a graph $G=(V,E)$, where $V$ represents the set of nodes (network components) and $E$ represents the set of edges (connections).
-Each node $v_{i}$ in the graph is associated with a feature vector $x_{i}$, which captures relevant information about the corresponding network component.
+The GNN can transform the graph structure data into standard representations, making them suitable for input into neural training.
+We employed the concept of class spectral filtering introduced in [49] and [50] to enhance the learning process.
+This approach allows the GNN to propagate the information efficiently from the nodes and edges to their neighboring nodes.
+We use the neighbor aggregation approach in [51] to achieve this propagation.
+The network infrastructure as depicted in Figure 3 is modeled as a graph $G=(V,E)$, where $V$ represents the node set (AMF, UDM, AUSF) and $E$ represents edge set.
+Each node $v_{i}$ in the graph is associated with a feature vector $x_{i}$, capturing relevant information regarding the corresponding network component.
 The graph can be represented using an adjacency matrix $A$, where $A_{ij}=1$ if a connection exists between nodes $v_{i}$ and $v_{j}$, and $A_{ij}=0$ otherwise.
-Node features are stored in the feature matrix $X$, where $X$ has the dimensions $|V|\times D$, with $|V|$ represents the number of nodes, and $D$ denotes the dimensionality of the feature vectors $X=[x_{1},x_{2},\ldots,x_{|V|}]$.
+Node features are stored in the feature matrix $X$, where $X$ has the dimensions $|V|\times D$, where $|V|$ represents the number of nodes, and $D$ denotes the dimensionality of the feature vectors $X=[x_{1},x_{2},\ldots,x_{|V|}]$.
 Many network components depend upon one another due to their interconnectivity [52].
-The GNN architecture consists of multiple layers, each processing information from neighboring vertices to extract hierarchical representations.
+The GNN architecture comprises of multiple layers, each processing information from neighboring vertices to extract hierarchical representations.
 We let $h_{i}^{(l)}$ denote the representation of vertex $v_{i}$ at layer $l$ of the GNN.
 Information is aggregated from the neighboring vertices and transformed using learnable parameters.
-The update equation for $h_{i}^{(l)}$ can be expressed as follow:
+The update equation for $h_{i}^{(l)}$ can be expressed as follows:
 $h_{i}^{(l)}=f^{(l)}\left(h_{i}^{(l-1)},\{h_{j}^{(l-1)}\}_{j\in N(v_{i})}\right),$ (1)
 where $N(v_{i})$ represents the set of neighboring vertices of $v_{i}$.
-One of the biggest challenges is creating network digital twin datasets that accurately replicate the real-world failure scenario.
-This difficulty arises to several factors, such as the limited availability of real-world system data and the inherent complexity of simulating diverse failure types.
+One of the greatest challenges is creating network digital twin datasets that accurately replicate the real-world failure scenarios.
+This problem arises from several difficulties, including the limited availability of real-world system data and the inherent complexity of simulating various types of failure.
 2.2 Graph Fourier transform
-We considered an undirected weighted graph $G=(V,E,w)$, where $V=\{0,1,\ldots,N-1\}$ is the vertex set, $E$ is the set of edges, and $w(i,j)$ is the weight function satisfying $w(i,j)=0$ for any $(i,j)$notin $E$.
-We assume all graphs are simple, meaning they have no loops or multiple edges.
-Three matrices associated with $G$ are important: the adjacency matrix $W$, a degree matrix $D$, and a Laplacian matrix $L=D-W$. The Laplacian matrix is essential for Graph Fourier Transforms (GFTs).
+This work considers an undirected weighted graph $G=(V,E,w)$, where $V=\{0,1,\ldots,N-1\}$ denotes the vertex set, $E$ represents the edge set, and $w(i,j)$ indicate the weight function satisfying $w(i,j)=0$ for any $(i,j)$notin $E$.
+This graph is assumed to be simple, meaning it contains no self-loops.
+Each node in this graph corresponds to a functional entity in the 5G core architecture, and edges represent measurable interactions (e.g., packet transfer, registration signaling, GTP flows).
+These matrices associated with $G$ are important: the adjacency matrix $W$, a degree matrix $D$, and a Laplacian matrix $L=D-W$.
+The Laplacian matrix is essential for GFTs.
 The Laplacian matrix $L$ is real, symmetric, and positive-semidefinite, thus, it possesses nonnegative eigenvalues $\lambda_{0},\ldots,\lambda_{N-1}$ and the corresponding orthonormal eigenfunctions $u_{0},\ldots,u_{N-1}$.
@@
 $L\begin{bmatrix}u_{k}(0)\\ \vdots\\ u_{k}(N-1)\end{bmatrix}=\lambda_{k}\begin{bmatrix}u_{k}(0)\\ \vdots\\ u_{k}(N-1)\end{bmatrix}$ (2)
-For $k=0,\ldots,N-1$, orthonormality implies that the sum of products of corresponding eigenfunctions is equal to the Kronecker delta function $\delta(i,j)$, where $\delta(i,j)$ equals 1 if $i=j$ and is zero otherwise.
+For $k=0,\ldots,N-1$, orthonormality implies that the sum of the products of the corresponding eigenfunctions equal the Kronecker delta function $\delta(i,j)$, where $\delta(i,j)$ equals 1 if $i=j$ and is zero otherwise.
 This study assumes that eigenvalues are arranged in ascending order, such that $\lambda_{0}\leq\ldots\leq\lambda_{N-1}$.
 In addition, $\lambda_{0}$ is strictly zero because the sum of rows in $L$ equals zero.
-The spectrum of the matrix, denoted as $\{\lambda_{k}\}_{k=0}^{N-1}$, is represented as $\sigma(L)$.
+The matrix spectrum, denoted by $\{\lambda_{k}\}_{k=0}^{N-1}$, is represented as $\sigma(L)$.
 The GFT in [53] of a graph feature $f:V\rightarrow\mathbb{R}$ is defined as $\hat{f}:\sigma(L)\rightarrow\mathbb{C}$, where $\sigma(L)$ represents the spectrum of the graph Laplacian $L$.
@@
 $\hat{f}(\lambda_{k})=\langle f,u_{k}\rangle=\sum_{i=0}^{N-1}f(i)u_{k}(i),$ (3)
-for $k=0,\ldots,N-1$. where, $u_{k}$ denotes the orthonormal eigenfunctions of $L$, and the GFT represents a feature expansion using these eigenfunctions. The inverse GFT is given by,
+for $k=0,\ldots,N-1$, where, $u_{k}$ denotes the orthonormal eigenfunctions of $L$, and the GFT represents a feature expansion using these eigenfunctions. The inverse GFT is given by,
 $f(i)=\sum_{k=0}^{N-1}\hat{f}(\lambda_{k})u_{k}(i),$ (4)
-which reconstruct the original function $f(i)$ from its spectral components.
-Interestingly, the GFT’s basic function in spectral graph research is highlighted by the fact that it is equal to the Discrete Fourier Transform (DFT) on cycle graphs.
+which reconstructs the original function $f(i)$ from its spectral components.
+The basic function of the GFT in spectral graph research is highlighted because it is equal to the discrete Fourier transform on cycle graphs.
 When a graph Laplacian has non-distinct eigenvalues, the functions generated by the GFT may not be well-defined, resulting in multi-valued functions.
@@
 $w((i_{1},i_{2}),(j_{1},j_{2}))=w_{1}(i_{1},j_{1})\delta(i_{2},j_{2})+\delta(i_{1},j_{1})w_{2}(i_{2},j_{2})$ (5)
-where, $\delta$ denotes the Kronecker delta function. The graphs $G_{1}$ and $G_{2}$ are referred to as the factor graphs of $G_{1}\square G_{2}$.
-It has been established in [59] that the adjacency, degree, and Laplacian matrices of Cartesian product graph can be derived form those of its factor graphs.
+where $\delta$ denotes the Kronecker delta function. The graphs $G_{1}$ and $G_{2}$ are referred to as the factor graphs of $G_{1}\square G_{2}$.
+The adjacency, degree, and Laplacian matrices of the Cartesian product graph can be derived from those of its factor graphs [54].
 Two factor graphs, $G_{1}$ and $G_{2}$, with vertex sets $V_{1}=\{0,1,\ldots,N_{1}-1\}$ and $V_{2}=\{0,1,\ldots,N_{2}-1\}$ respectively are considered.
-Each with adjacency matrix $W_{1}$ and $W_{2}$, degree matrix $D_{1}$ and $D_{2}$, and Laplacian matrix $L_{1}$ and $L_{2}$.
-When the vertices of the Cartesian product graph are ordered lexicographically, such as $(0,0),(0,1),(0,2),\ldots,(N_{1}-1,N_{2}-1)$, the adjacency, degree, and Laplacian matrices of $G_{1}\square G_{2}$ can be expressed as $W_{1}\oplus W_{2}$, $D_{1}\oplus D_{2}$, and $L_{1}\oplus L_{2}$, respectively, where operator $\oplus$ denotes the Kronecker sum.
-Definition 1: The Kronecker sum is defined by $A\oplus B=A\otimes I_{n}+I_{m}\otimes B$ for matrices $A\in\mathbb{R}^{m\times m}$ and $B\in\mathbb{R}^{n\times n}$, where $I_{n}$ represents the identity matrix of size $n$.
+Each with adjacency matrix $W_{1}$ and $W_{2}$, degree matrix $D_{1}$ and $D_{2}$, and Laplacian matrix $L_{1}$ and $L_{2}$ respectively.
+When the vertices of the Cartesian product graph are ordered lexicographically, such as $(0,0),(0,1),(0,2),\ldots,(N_{1}-1,N_{2}-1)$, the adjacency, degree, and Laplacian matrices of $G_{1}\square G_{2}$ can be expressed as $W_{1}\oplus W_{2}$, $D_{1}\oplus D_{2}$, and $L_{1}\oplus L_{2}$, respectively, where the operator $\oplus$ denotes the Kronecker sum.
+Definition 1: The Kronecker sum is defined by $A\oplus B=A\otimes I_{n}+I_{m}\otimes B$ for matrices $A\in\mathbb{R}^{m\times m}$ and $B\in\mathbb{R}^{n\times n}$, where $I_{n}$ represents an identity matrix of size $n$.
 The Kronecker sum has a valuable characteristic that decomposes an eigenproblem involving the Laplacian matrix of a product graph into eigenproblems of Laplacian matrices of the factor graphs.
@@
 The resulting eigenvalues $\{\lambda^{(1)}_{k}+\lambda^{(2)}_{k}\}$ are nonnegative, and the corresponding eigenfunctions $\{u^{(1)}_{k}\otimes u^{(2)}_{k}\}$ are orthonormal.
-These conclusions can be straightforwardly derived from the fundamental properties of the Kronecker product.
+These conclusions can be directly derived from the fundamental properties of the Kronecker product.
 A similar approach applies to decomposing an eigenproblem related to an adjacency matrix of a product graph.
 If the adjacency matrix $W_{n}$ has eigenvalues $\{\mu^{(n)}_{k}\}_{k=0}^{N_{n}-1}$ and orthonormal eigenfunctions $\{\mathbf{v}^{(n)}_{k}\}_{k=0}^{N_{n}-1}$ for $n=1,2$, the Kronecker sum $W_{1}\oplus W_{2}$ yields the eigenvalues $\{\mu^{(1)}_{k}+\mu^{(2)}_{l}\}$ and corresponding eigenfunctions $\{\mathbf{v}^{(1)}_{k}\otimes\mathbf{v}^{(2)}_{l}\}$ defined on $V_{1}\times V_{2}$.
-The Cartesian product graph is represented as $G_{1}\times G_{2}$, where $G_{n}$ for $n=1,2$ is an undirected weighted graph with a vertex set $V_{n}=\{0,1,\dots,N_{n}-1\}$.
-Assuming its graph Laplacian $L_{n}$ has eigenvalues $\{\lambda_{k}^{(n)}\}_{k=0}^{N_{n}-1}$ and corresponding orthonormal eigenfunctions $\{u_{k}^{(n)}\}_{k=0}^{N_{n}-1}$.
-The GFT of a graph feature $f:V_{1}\times V_{2}\rightarrow\mathbb{R}$ on the product graph $G_{1}\times G_{2}$ is represented by:
+The Cartesian product graph is represented as $G_{1}\times G_{2}$, where $G_{n}$ for $n=1,2,G_{n}$ is an undirected weighted graph with a vertex set $V_{n}=\{0,1,\dots,N_{n}-1\}$.
+We assume its graph Laplacian $L_{n}$ has eigenvalues $\{\lambda_{k}^{(n)}\}_{k=0}^{N_{n}-1}$ and corresponding orthonormal eigenfunctions $\{u_{k}^{(n)}\}_{k=0}^{N_{n}-1}$.
+The GFT of a graph feature $f:V_{1}\times V_{2}\rightarrow\mathbb{R}$ on the product graph $G_{1}\times G_{2}$ is represented by,
 $\hat{f}:(\sigma(L_{1})\otimes\sigma(L_{2}))\rightarrow\mathbb{C}$ (16)
@@
 $f(i_{1},i_{2})=\sum_{k_{1}=0}^{N_{1}-1}\sum_{k_{2}=0}^{N_{2}-1}\hat{f}(\lambda^{(1)}_{k_{1}}+\lambda^{(2)}_{k_{2}})\cdot u^{(1)}_{k_{1}}(i_{1})\cdot u^{(2)}_{k_{2}}(i_{2})$ (18)
-Thus, we interpret features on a Cartesian product graph as 2D features and propose a Twin GFT, enabling multiclass failure classification in Network Digital Twin environments.
-Definition 1 The Twin Graph Fourier Transform (Twin-GFT) of a function $f:V_{1}\times V_{2}\rightarrow\mathbb{R}$ on a Cartesian product graph $G_{1}\square G_{2}$ is a spectral representation $\hat{f}:\sigma(L_{1})\times\sigma(L_{2})\rightarrow\mathbb{C}$ defined as follows:
+Thus, this work interprets features on a Cartesian product graph as two-dimensional features and proposes a twin GFT, enabling multiclass failure classification in Network Digital Twin environments.
+Definition 1 The twin GFT of a function $f:V_{1}\times V_{2}\rightarrow\mathbb{R}$ on a Cartesian product graph $G_{1}\square G_{2}$ is a spectral representation $\hat{f}:\sigma(L_{1})\times\sigma(L_{2})\rightarrow\mathbb{C}$ defined as follows:
 $\hat{f}(\lambda_{11},\lambda_{22})=\sum_{i_{1}=0}^{N_{1}-1}\sum_{i_{2}=0}^{N_{2}-1}f(i_{1},i_{2})\cdot u_{k_{11}}^{(1)}(i_{1})\cdot u_{k_{22}}^{(2)}(i_{2})$ (19)
@@
 for $i_{1}=0,\dots,N_{1}-1$ and $i_{2}=0,\dots,N_{2}-1$.
-The twin-GFT can be represented as a series of matrix-matrix multiplications.
-Using $N_{1}\times N_{2}$ matrices $F=(f(i_{1},i_{2}))_{i_{1},i_{2}}$ and $\hat{F}=(\hat{f}(\lambda_{k_{11}}^{(1)}+\lambda_{k_{22}}^{(2)}))_{k_{11},k_{22}}$, the twin-GFT applied to the features $f$ is expressed as follows:
+The twin GFT can be represented as a series of matrix-to-matrix multiplications.
+The twin GFT applied to features f is expressed using $N_{1}\times N_{2}$ matrices $F=(f(i_{1},i_{2}))_{i_{1},i_{2}}$ and $\hat{F}=(\hat{f}(\lambda_{k_{11}}^{(1)}+\lambda_{k_{22}}^{(2)}))_{k_{11},k_{22}}$, as follows:
 $\hat{F}=U_{1}^{*}FU_{2},$ (21)
-where $U_{n}$ denotes an $N_{n}\times N_{n}$ unitary matrix with $(i,k)$-th element $u_{k}^{(n)}(i)$ for $n=1,2$. Then its inverse transform is given by:
+where $U_{n}$ denotes an $N_{n}\times N_{n}$ unitary matrix with the $(i,k)$-th element $u_{k}^{(n)}(i)$ for $n=1,2$. Then, its inverse transform is given as follows:
 $F=U_{1}\hat{F}^{2}U_{2}^{*},$ (22)
 The twin GFT has connections to existing transformations.
-First, when both factor graphs are cycle graphs, the twin GFT can be equivalent to the 2D-Discrete Fourier Transform and the 2D-GFT.
+First, when both factor graphs are cycle graphs, the twin GFT can be equivalent to the 2D-Discrete Fourier Transform and the 2D-discrete Fourier transform and 2D GFT.
 Second, when a factor graph is a cycle graph, some cycles might be nested within other cycles, creating a complex network structure.
-The twin GFT is known as a joint graph and temporal Fourier transform which generalizes these existing transformations.
-3 PROPOSED CF-GNN
-Class-Fourier graph Neural Networks (CF-GNNs) perform well with both homophilic and heterophilic graphs and have shown good expressive capacities.
-However, current methods often lack adaptability.
-In particular, conventional approaches use a single, globally shared spectral filter $\hat{g}$, trained across the entire graph, with fixed frequency coefficients $\{\gamma_{k}\}_{k=0}^{K}$.
-Node-specific changes are not considered in the global filter, which is unable to distinguish between the distinct local structures connected to each node during the filtering process.
+The twin GFT is a joint graph and temporal Fourier transform, generalizing the existing transformations.
+3 Proposed Class-Fourier Graph Neural Network
+The CF-GNN performs well with homophilic and heterophilic graphs and has demonstrated good expressive capacities.
+However, the current methods often lack adaptability.
+Conventional approaches apply a single, globally shared spectral filter $\hat{g}$, trained across the entire graph, with fixed frequency coefficients $\{\gamma_{k}\}_{k=0}^{K}$.
+Node-specific changes are not considered in the global filter, which cannot distinguish between the distinct local structures connected to each node during the filtering process.
 Although polynomial-parameterized spectral filter learning offers a certain level of localization, it might not adequately capture subtle and diverse local structural patterns.
-It is intuitively reasonable to train a class-specific filter $\hat{g}_{i}(\lambda_{l})$ rather than depending on a globally consistent spectral filter $\hat{g}(\lambda_{l})$.
-A more efficient method is provided by $\hat{g}_{i}(\lambda_{l})$, tailored for each node $i$ to adaptively capture its local patterns.
-To mitigate this limitation, we revisit globally consistent spectral graph transforms in this section and propose a spectral filter learning strategy in the context of CF-GNNs.
+It is intuitively reasonable to train a class-specific filter $\hat{g}_{i}(\lambda_{l})$ rather than to depend on a globally consistent spectral filter $\hat{g}(\lambda_{l})$.
+A more efficient method is provided by $\hat{g}_{i}(\lambda_{l})$, tailored for each node $i$ to adaptively capture its local patterns adaptively.
+This section revisits globally consistent spectral graph transforms and proposes a spectral filter learning strategy in the context of CF-GNNs to mitigate this limitation.
 3.1 Class-Fourier transform filtering
-We present an adaptive localized spectral filtering method designed for Class-Fourier Transform Filtering on graph $G$, which the generalized translation operator inspires in Graph Signal Processing (GSP) in [60].
-By leveraging polynomial-parameterized spectral filtering from [61], this technique ensures a more accurate and flexible representation of local patterns by fully accounting for the unique influence of the node at which the filter is applied.
-We introduce a generalized translation operator $T_{i}$ for any signal $g\in\mathbb{R}^{n}$ defined on a given graph $G$ and any $i\in\{0,1,\dots,n-1\}$: $T_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n}$
-Using a Kronecker delta function in (17) and (18) $\delta_{i}$ centered at the $i$-th node $v_{i}$, we express generalized convolution as:
-$T_{i}(g):=\sqrt{N}\,(g\ast\delta_{i})=\sqrt{N}\sum_{l=1}^{n}u_{l}u_{l}^{\top}(i)\hat{g}(\lambda_{l})$ (23)
-where $u_{l}^{\top}(i)$ denotes the $i$-th element of the eigenvector $u_{l}$ corresponding to eigenvalue $\lambda_{l}$, and $\hat{g}(\lambda_{l})$ represents the spectral domain representation of the signal $g$.
-In the context of Class-Fourier Transform Filtering, adaptive local filtering is achieved by first centering the filter signal $g$ at the target node $v_{i}$ using the generalized translation operator $T_{i}$, and then performing spectral convolution with $x$ as follows:
+This section introduces an adaptive localized spectral filtering method for CFT filtering on a graph $G$, inspired by the generalized translation operator from graph signal processing (GSP) in [21].
+Using polynomial-parameterized spectral filtering as described in [55], this approach ensures accurate and flexible representation of local patterns by incorporating the influence of the node at which the filter is applied.
+We define a generalized translation operator $T_{i}$ for any signal $g\in\mathbb{R}^{n}$ defined on a graph $G$, and for any node $i\in\{0,1,\dots,n-1\}$, as:
+$T_{i}:\mathbb{R}^{n}\to\mathbb{R}^{n},$ (23)
+and specifically,
+$T_{i}(g):=\sqrt{N}\,(g\ast\delta_{i})=\sqrt{N}\sum_{l=1}^{n}u_{l}u_{l}^{\top}(i)\hat{g}(\lambda_{l}),$ (24)
+where $u_{l}^{\top}(i)$ denotes the $i$-th entry of the eigenvector $u_{l}$ corresponding to eigenvalue $\lambda_{l}$, and $\hat{g}(\lambda_{l})$ is the spectral representation of signal $g$.
+In the context of CFT filtering, adaptive local filtering is achieved by centering the filter signal $g$ at the target node $v_{i}$ via $T_{i}$, followed by spectral convolution with $x$:
 $x\ast GT_{i}(g)=\sum_{l=1}^{N}u_{l}\,x^{(\lambda_{l})}u_{l}^{\top}(i)g^{(\lambda_{l})}$ (25)
-where $x^{(\lambda_{l})}$ is the spectral domain representation of the input signal $x$, and $u_{l}^{\top}(i)$ is the $i$-th element of the eigenvector $u_{l}$, which corresponds to the eigenvalue $\lambda_{l}$.
-Establishing that: $\hat{g}_{i}(\lambda_{l})=\sqrt{N}\,u_{l}(i)\,\hat{g}(\lambda_{l})$ in 24, the equation above becomes:
+Since $\hat{g}_{i}(\lambda_{l})=\sqrt{N}\,u_{l}(i)\,\hat{g}(\lambda_{l})$, we can rewrite Equation 25 as:
 $x\ast GT_{i}(g)=\sum_{l=1}^{N}u_{l}\,x^{(\lambda_{l})}g_{i}^{(\lambda_{l})}$ (26)
-As evident from the definition of $\hat{g}_{i}(\lambda_{l})$, $\hat{g}_{i}$ is a function associated with $U$.
-Thus, by approximating $U$, we can further establish the relationship between $\hat{g}_{i}$ and $x$.
-Recall the Fourier transform of an inverted graph, which allows us to deduce that $x_{i}=U_{i:}\cdot\hat{x},$ where $U_{i:}$ denotes $U$’s $i$-th row.
-Therefore, $U_{i:}$ can be approximately from $x_{i}$ using $U_{i:}\approx x_{i}q,$where $\hat{x}$’s pseudoinverse is $q=\text{pinv}(\hat{x})\in\mathbb{R}^{n}$.
-Meanwhile, take notice that $u_{l}(i)=\tilde{u}_{l}(i)$, which implies that $\hat{g}_{i}(\lambda_{l})$ can be approximately expressed as $\hat{g}_{i}(\lambda_{l})\approx\ldots$.
-$\hat{g}_{i}(\lambda_{l})\approx\tilde{g}_{i}(\lambda_{l})=\sqrt{N}(x_{i}q_{l})\hat{g}(\lambda_{l})$ (26)
-Where the $q_{l}$ denotes the $l$-th element of $\mathbf{q}$.
-Consequently, a particular filter $\tilde{g}_{i}$ corresponding to node $v_{i}$, can approximate $\hat{g}_{i}$ while simultaneously incorporating the influence of feature $x_{i}$ associated with to node $v_{i}$ in the filter coefficient estimation process.
-Without loss of generality, the $K$-order polynomial approximation $\hat{P}_{K}(\cdot)$ can be utilized to directly parameterize $\tilde{g}_{i}$ without sacrificing generality; that is,$\tilde{g}_{i}=\sum_{k=0}^{K}\eta_{i,k}\hat{p}_{k}(\Lambda),$ where $\eta_{i,k}$ represents the coefficient of $\hat{p}_{k}(\Lambda)$ for the $\tilde{g}_{i}$ filter.
-We only pay attention to the convolution value of the filter placed at each node is our focus, i.e.,$z_{i}=\delta_{i}(U\tilde{g}_{i}U^{\top}x).$ In this case, we utilize $\delta_{i}$ as an indicator vector $\delta_{i}=[0,\ldots,1_{i},\ldots,0]\in\mathbb{R}^{n},$ which represents a row vector with only the $i$-th element being 1 and the remaining elements being zeros.
-Therefore, the following is a formulation of class-oriented filtering:
-$z_{i}=\delta_{i}\big{(}U\tilde{g}_{i}U^{\top}x\big{)}=\delta_{i}U\sum_{k=0}^{K}\eta_{i,k}\hat{p}_{k}(\Lambda)U^{\top}x$
+Noting that $x_{i}=U_{i:}\cdot\hat{x}$, where $U_{i:}$ is the $i$-th row of the eigenvector matrix $U$, and $q=\text{pinv}(\hat{x})\in\mathbb{R}^{n}$, we approximate:
+$U_{i:}\approx x_{i}q,\quad\hat{g}_{i}(\lambda_{l})\approx\tilde{g}_{i}(\lambda_{l})=\sqrt{N}(x_{i}q_{l})\hat{g}(\lambda_{l})$ (27)
+with $q_{l}$ the $l$-th element of $q$.
+This leads to a node-specific filter $\tilde{g}_{i}$ that reflects the influence of feature $x_{i}$. The filter can be parameterized using a $K$-order polynomial approximation:
+$\tilde{g}_{i}=\sum_{k=0}^{K}\eta_{i,k}\hat{p}_{k}(\Lambda),$ (28)
+where $\eta_{i,k}$ are trainable coefficients. Focusing on the filter output at each node, we compute:
+$z_{i}=\delta_{i}(U\tilde{g}_{i}U^{\top}x)=\delta_{i}U\sum_{k=0}^{K}\eta_{i,k}\hat{p}_{k}(\Lambda)U^{\top}x$ (29)
 $=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\hat{p}_{k}(L)x$ (30)
-Where the trainable coefficient matrix is $\Psi=[\eta_{i,k}]\in\mathbb{R}^{n\times(K+1)}$ In practice,
+with $\Psi=[\eta_{i,k}]\in\mathbb{R}^{n\times(K+1)}$. This extends to a feature matrix $X$ as:
 $Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\hat{p}_{k}(L)X$ (31)
-Eq (28) can still be applied to the feature matrix $X$.
+CF-GNN-V: Eigenvalue-based Transformation
+This variant applies a transformation to the learned features in the spectral domain that emphasizes eigenvalue-related frequency components relevant to the minority class.
+It introduces a class-sensitive weighting function $\gamma(\lambda)$ such that:
+$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\gamma(\Lambda)\hat{p}_{k}(L)X$ (32)
+Here, $\gamma(\Lambda)$ is a diagonal matrix scaling frequency components to highlight class-specific spectral patterns.
+Proof of CF-GNNv
+The base form of graph convolution in the spectral domain is:
+$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\hat{p}_{k}(L)X$ (33)
+This corresponds to filtering the signal $X$ using polynomial spectral filters $\hat{p}_{k}(\Lambda)$, followed by node-wise localization using $\delta_{i}$.
+In CF-GNNv, we introduce a class-sensitive transformation $\gamma(\Lambda)$ into the filtering process:
+$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\gamma(\Lambda)\hat{p}_{k}(L)X$ (34)
+In spectral graph theory, filtering a signal $x\in\mathbb{R}^{n}$ is represented as:
+$x^{\prime}=Ug(\Lambda)U^{\top}x$ (35)
+where $g(\Lambda)$ is a diagonal matrix that applies the filter in the spectral domain.
+To emphasize frequency components relevant to the minority class, we define $\gamma(\lambda)$ to selectively modulate these components.
+Since $\gamma(\Lambda)$ and $\hat{p}_{k}(\Lambda)$ are diagonal, we can define a composite filter:
+$\tilde{g}_{i,k}(\Lambda)=\gamma(\Lambda)\hat{p}_{k}(\Lambda)$ (36)
+which leads to:
+$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}U\tilde{g}_{i,k}(\Lambda)U^{\top}X$
+$=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}U\gamma(\Lambda)\hat{p}_{k}(\Lambda)U^{\top}X$
+$=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\gamma(\Lambda)\hat{p}_{k}(L)X$ (37)
+This final expression shows that CF-GNN-V modifies the spectral filter to reflect class-sensitive frequency emphasis.
+3.1.1 Eigenvector-aware attention
+We begin with the classical spectral graph convolution expression, as shown earlier in Eq. (35):
+$x^{\prime}=Ug(\Lambda)U^{\top}x$ (38)
+Now, instead of only manipulating the eigenvalues $\Lambda$ using $g(\Lambda)$, we propose to modify the spectral projection basis $U$ using eigenvector-aware attention $\alpha(U)$, where $\alpha(U)\in\mathbb{R}^{n\times n}$ could be either a full matrix or a diagonal one (simpler to implement and more stable in training).
+The modified filtering of the signal is:
+$x^{\prime}=U\alpha(U)g(\Lambda)U^{\top}x$ (39)
+This formulation applies an attention mechanism in the spectral domain on the eigenvectors, rather than just adjusting the spectral coefficients (eigenvalues).
+The idea is that certain eigenvectors capture structural features (e.g., localized clusters, boundary conditions) that may be more informative for classifying failure modes, especially in minority classes.
+We define the composite filter with eigenvector attention as:
+$\tilde{g}_{i,k}(U,\Lambda)=\alpha(U)\hat{p}_{k}(\Lambda)$ (40)
+Substituting into the spectral convolution yields:
+$Z_{i}=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}U\alpha(U)\hat{p}_{k}(\Lambda)U^{\top}X=\delta_{i}\sum_{k=0}^{K}\Psi_{i,k}\alpha(U)\hat{p}_{k}(L)X$ (41)
+This final expression shows that CF-GNNE introduces class-aware attention over eigenvectors in the spectral filtering process, complementing the eigenvalue-focused approach of CF-GNNv.
 Input: Symmetric matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$
@@
 Ensemble classification result $H(v)\leftarrow$ from all weak classifiers;
-Algorithm 1 CF-GNN: Class Fourier Graph Neural Network
+Algorithm 1 CF-GNN: Class-Fourier Graph Neural Network
 3.2 Experimental settings
-3.3 Datasets and preprocessing
-1) KDDI datasets: The KDDI dataset is based on the ITU Challenge ”ITU-ML5G-PS-008: Network Failure Classification Model Using network digital twin” [45].
-The datasets consist of the following: a dataset (Domain A) derived from network digital twin simulated environments and a dataset (Domain C) generated from a real network.
-Each domain dataset comprises 4121 features, with 16 failure classes per failure sample.
-However, the failure-type classes consist of a ”normal” class and 15 other classes, with over 67% of the data belonging to the ”normal” class.
-In comparison, each of the remaining 15 failure classes comprises only 2.2% of the datasets.
-Table 2 represents the distribution table for the two datasets
-2) Pre-processing: This section details the data preprocessing stage of the system.
-This stage consists of two subprocesses: one-hot encoding for categorical classes, and feature scaling for normalization.
-The system transforms nominal attributes into one-hot vectors.
-Each nominal (categorical) attribute is represented as a binary vector with a size equal to the number of attribute values.
-In this binary vector, only one point corresponds to the expressed value, which is assigned a value of 1, whereas all other points are assigned a value of 0.
-For instance, for the failure injection attribute commonly used in network failure classification, with the failure points pf the process and mobility management function “amf,” authentication server function “ausf,” and unified data management “udm,” the attribute is transformed into a binary vector of length 3.
-The attribute values are converted to [1, 0, 0], [0, 1, 0], and [0, 0, 1], respectively.
-The system scales the numeric attributes alongside the one-hot encoding process.
-Normalization methods, such as normalization (e.g., [62]) and standardization (e.g., [63]) can be considered for scaling numeric features.
-We adopted the min-max normalization method.
-The normalization function $\mathcal{f}_{A}(\cdot)$ for a numeric attribute $A$ that maps every $x$ in $A$ into the range [0, 1] can be defined as follows:
-$f_{A}(x_{i})=\tilde{x}_{i}=\frac{\text{max}(x_{j})-\text{min}(x_{j})}{x_{i}-\text{min}(x_{j})},$ (29)
-where, $x_{i}$ denotes the $i$th attribute value in attribute $A$.
+3.2.1 Description of the KDDI datasets
+The KDDI dataset is based on the ITU Challenge ”ITU-ML5G-PS-008: Network Failure Classification Model Using Network Digital Twin” [38].
+The dataset comprises (Domain A) derived from the simulated environment of a network digital twin, and (Domain C) a dataset generated from a real-world network.
+Each domain dataset comprises 4,121 features, with 16 failure classes per failure sample.
+However, failure-type classes consist of a “normal” class and 15 other classes, with more than 67% of the data belonging to the “normal” class.
+In comparison, the remaining 15 failure classes comprise only 2.2% of the dataset.
+Table 3 presents the class distribution for the two datasets, whereas Table 2 lists each failure point along with its corresponding failure types, some features, and descriptions relevant to the datasets used.
+Table 2: Description of the dataset based on failure nodes, failure points, classes, and some features used for the classification.
+Node Types Classes Features Description
+AMF amfx1 bridge-delif amfx1_statistics.out-octets Tracks AMF packet flow for mobility. Monitors CPU/memory load to identify failure.
+ens5 inter-down out-unicast-pkts
+ens5 interface-loss memory-stats.used-percent_value
+memory-stress-start per-core-CPU-system_value
+vcpu-overload-start
+AUSF ausfx1 bridge-delif RM.RegisteredSubNbrMax Measures registered subscribers for authentication. Monitors GTP traffic from AUSF
+inter-down RM.RegisteredSubNbrMean
+inter-loss-70 GTP.OutDataOctN3UPF
+memory-stress-start GTP.OutDataPktN3UPF
+vcpu-overload-start
+UDM udmx1 bridge-delif udmx1_five-g.RM.RegInitReq Handles user profile access and registration. Tracks traffic for UDM data management.
+inter-down udmx1_ens5_statistics.out-unicast-pkts_value
+inter-loss-70 udmx1_ens5_speed_value
+memory-stress-start ens5.out-unicast-pkts
+vcpu-overload-start ens5.speed
+3.2.2 Parameter settings
+We randomly select the training samples while maintaining the same positive-to-negative sample ratio as in the entire dataset.
+The Adam optimizer is employed for parameter optimization, and the hyperparameters are configured as follows: For the network digital twin dataset (Domain A), the hidden embedding size is set to 64 with a $L_{2}$ regularization value of $5e{-4}$, the learning rate to 0.01, the adjacency dropout to 0.2, with two model layers, and training is conducted for 350 epochs.
+For the real-world network dataset (Domain C), the hidden embedding size is also set to 64, weight loss $(\lambda)$ is 1e-6, learning rate to 0.01, the number of model layers to 2, and the maximum number of epochs to 250, and imbalance ratio is set 0.1 to 0.9.
+The proposed method is implemented using PyTorch and is executed in a Python 3.8 environment on a Windows 10 machine equipped with a 64-bit operating system, a 13th-generation Intel Core i5-13400 processor, 32 GB of RAM, and an NVIDIA GeForce RTX 3060 Ti GPU.
+3.2.3 Preprocessing
+To prepare the data for input into the GNN model, we performed two primary preprocessing steps: categorical encoding and feature normalization.
+Categorical attributes are transformed into binary vectors via one-hot encoding.
+Given a categorical attribute with $k$ possible values, we create a binary vector of length $k$, where only the index corresponding to the actual category is set to 1.
+For instance, Figure 3 represents the components of the three failure point functions (AMF, AUSF, and UDM) and is encoded as $[1,0,0]$, $[0,1,0]$, or $[0,0,1]$, respectively.
+To ensure scale invariance and stabilize training, we apply min-max normalization [56] to all numeric features. The transformation is defined by:
+$f_{A}(x_{i})=\tilde{x}_{i}=\frac{x_{i}-\min(x_{j})}{\max(x_{j})-\min(x_{j})},$ (42)
+where $x_{i}$ is the $i^{\text{th}}$ value of the numeric attribute $A$, and $\min(x_{j})$, $\max(x_{j})$ denote the minimum and maximum values in attribute $A$, respectively.
+This mapping scales all values into the range $[0,1]$, which improves convergence and comparability between different types of features.
 Table 3: Dataset distribution table
 Dataset Failure events Features IR
-Domain A (Training) Normal = 2456 4121 0.4886
+Domain A Normal = 2456 4,121 0.4886
 Failure = 1186
-Domain C (Training) Normal = 588 4121 0.4847
+Domain C Normal = 588 4,121 0.4847
 Failure = 285
 3.3 Comparison with the baseline methods
-The CF-GNN presented in this article is fundamentally a GNN model designed to identify failure classes in a network digital twin setting.
+The CF-GNN presented in this article is a GNN model designed to identify failure classes in a network digital twin setting.
 The primary scientific problem addressed is graph representation learning for imbalanced graphs.
-We compared our results with cutting-edge baseline techniques that include: 1) general GNN models GCN [41] and GAT [15]; 2) Four imbalance-based models, including PC-GNN [22], Reweight [54], GraphSMOTE [27], and GATE-GNN [29] as state-of-art GNN imbalance classification models.
+We compared the results with innovative baseline techniques that include: 1) general GNN models, GCN [34], and GAT [57]; and 2) four imbalance-based models, including PC-GNN [8], Reweight [48], GraphSMOTE [13], and GATE-GNN [15], as state-of-art GNN imbalance classification models.
   • 1.
-    GCN: GCN aggregates information and updates its presentation of data from the node’s neighbors, repeating this process across several layers.
+    GCN: The GCN aggregates information and updates its data presentation from the neighboring nodes, repeating this process across several layers.
   • 2.
-    GAT: GAT offers more accurate and nuanced representations of the graph data compared to earlier GNN by enabling the network to learn which neighbors are most significant for each node.
+    GAT: The GAT offers more accurate representations of the graph than earlier GNNs by enabling the network to learn which neighbors are the most significant for each node.
   • 3.
-    Rewight: The core idea of gradient reweighting is to mitigate the effects of class imbalance by adjusting the gradient updates during training.
+    In Reweight Loss the core idea of gradient reweighting is to mitigate the effects of class imbalance by adjusting the gradient updates during training.
   • 4.
-    GraphSMOTE: GraphSMOTE is a technique that generates synthetic nodes while preserving the graph structure to address the challenge of imbalanced node classification in graphs.
+    GraphSMOTE technique generates synthetic nodes while preserving the graph structure to address the challenge of imbalanced node classification in graphs.
   • 5.
-    PC-GNN: The model is exposed to a more balanced training input by sampling balanced sub-graphs, which keeps the majority classes dominating the learning process.
+    PC-GNN is exposed to a more balanced training input by sampling balanced subgraphs, maintaining the domination of the majority classes in the learning process.
   • 6.
-    GATE-GNN: GATE-GNN model for imbalanced node classification. It creates dynamic node interaction within the GATE architecture using learnable weight parameters.
+    GATE-GNN method for a classification of imbalanced nodes, creates dynamic node interactions within the GATE architecture using learnable weight parameters.
   • 7.
-    CF-GNN: Our proposed CF-GNN model improves performance, particularly for minority classes in imbalanced graph classification, by leveraging the Fourier transform to examine class-specific frequency features.
-Evaluation Metrics: The selection of evaluation metrics is essential when addressing imbalanced datasets.
-These metrics must not only provide a fair assessment of classification results across all classes but also pay special attention to the performance of minority classes [64].
-To ensure an objective evaluation, we have chosen five widely used metrics: Precision, Recall, F1-score, G-Mean, and MCC.
+    In the CF-GNN, the proposed model improves performance, particularly for minority classes in imbalanced graph classification, by employing the Fourier transform to examine class-specific frequency features.
+  • 8.
+    CF-GNN_v applies a spectral-inspired transformation to amplify frequency components associated with class distinctions.
+  • 9.
+    CF-GNN_E: Uses attention over spectral embeddings to focus on structurally informative nodes. Boosts sensitivity to underrepresented class patterns in imbalanced graph data.
+3.4 Evaluation Metrics
+The selection of evaluation metrics is essential when addressing imbalanced datasets.
+These metrics must provide a fair assessment of classification results across all classes and pay special attention to the performance of minority classes [7].
+We chose five widely used metrics to ensure an objective evaluation: the precision, recall, F1-score, geometric mean, and Mathews correlation coefficient (MCC):
 $\text{Accuracy}=\frac{TP+TN}{TP+FP+TN+FN}$ (43)
@@
 where TP, TN, FN, and FP denote the true positive, true negative, false negative, and false positive values, respectively.
-In the CF-GNN experiments, we evaluated each model using standard performance metrics.
-The precision, recall, and F1-score metrics vary significantly across classes, indicating differences in how well the model predicts each class.
+The CF-GNN experiments, we evaluated each model using standard performance metrics.
+The precision, recall, and F1-score metrics vary significantly by class, indicating differences in how well the model predicts each class.
 This heterogeneity in performance suggests that the model may perform well for some classes and less effectively for others.
-Geometric-Mean (G-Mean) is computed using a True Positive Rate (TPR) and True Negative Rate (TNR).
-[…]
-A higher score for the aforementioned criteria indicates better model performance on imbalanced problems.
-The Matthews Correlation Coefficient (MCC) generalizes the confusion matrix for multiclass classification. The generalized equation is:
+The geometric mean is computed using a TP rate (TPR) and a TN rate (TNR).
+Geometric $\text{Geometric mean}=\sqrt{\text{TPR}\cdot\text{TNR}}=\sqrt{\frac{\text{TP}}{\text{TP}+\text{FN}}\cdot\frac{\text{TN}}{\text{TN}+\text{FP}}}.$ (47)
+A higher score for the mentioned criteria indicates better model performance on imbalanced problems.
+The MCC generalizes the confusion matrix for multiclass classification.
+The generalized equation is:
 $\text{MCC}=\frac{\sum_{k}\sum_{l}\sum_{m}C_{kk}C_{lm}-C_{kl}C_{mk}}{\sqrt{\left(\sum_{k}\left(\sum_{l}C_{kl}\right)\left(\sum_{l\neq k}C_{lk}\right)\right)\cdot\left(\sum_{k}\left(\sum_{l}C_{lk}\right)\left(\sum_{l\neq k}C_{kl}\right)\right)}}$
-where $k,l,m$ are indices corresponding to certain classes and $C_{kl}$ are elements of the confusion matrix.
+where $k,l,andm$ represent indices corresponding to certain classes and $C_{kl}$ denotes the elements of the confusion matrix.
 The MCC is particularly suitable for multiclass imbalanced situations because it ensures a balanced assessment across all classes.
@@
 This section provides an empirical analysis of the CF-GNN using two 5G network digital twin datasets.
-It begins with a description of the experimental setup, including the datasets and evaluation metrics, followed by a detailed evaluation and analysis of the experimental results.
-Table 3: Training Parameters
-Parameters Value
-Dimension Hidden_dim=64
-Learning rate lr = 0.001
-Optimizer Adam optimizer
-Loss function Cross-entropy loss
-Number of epochs num_epochs = 500
-Dimension Hidden_dim
+This section describes and analyzes the experimental results, including the datasets and evaluation metrics, followed by a detailed evaluation and discussion of the experimental results.
 4.1 Overall performance comparison
-In this section, we evaluate the effectiveness of the proposed CF-GNN technique, its performance was compared with the state-of-the-art models using the class balanced accuracy metric (cmA) for handling imbalanced classification tasks.
-The results shown in Figure 2 and 3 demonstrate that conventional models such as GCN and GAT, struggled with imbalanced datasets, often plateauing at lower accuracy levels.
-For example, GAT exihibits fluctuations during training, indicating instability, while GCN reaches an early performance maximum and fails to adapt effectively to minority classes.
-In contrast, other methods like PC-GNN and GraphSMOTE achieve better performance due to their enhanced node representations and oversampling approaches.
-However, techniques such as Reweight show significant instability, particularly in the initial stages of training, suggesting that merely adjusting loss functions is insufficient to address class imbalance effectively.
-Although GATE-GNN, which aggregates node features using adaptive weights, performs competitively, it still exhibits some inconsistencies during the training.
-Figure 2: Classification Mean Accuracy (cmA) from (Domain C) models performance comparison
-Figure 3: Classification Mean Accuracy (cmA) from (Domain C) models performance comparison
-On the other hand, CF-GNN (ours) consistently outperforms all other models across every dataset.
-It avoids the erratic fluctuations observed in Reweight and other baselines, achieving a smoother and more stable convergence.
-The main advantages of CF-GNN lies in its class-oriented filtering mechanism, where the algorithm iterates over classes to ensure the filter learns representations for each class.
-Moreover, this method ensures that the spectral filtering process focuses on class-dependent graph structure and features, leading to better generalization across tasks and improved classification accuracy, particularly for minority classes.
-The performance results demonstrate that CF-GNN outperforms the ebst baseline models by 1-3% in cmA, highlighting its effectiveness in addressing the challenges posed by long-tail distributions.
-This enhancement positions CF-GNN as a reliable and scalable solution for practical network digital twin applications, underscoring the importance of integrating spectral approaches with graph-based imbalance classification.
+This section evaluates the effectiveness of the proposed CF-GNN technique, and its performance was compared with state-of-the-art models using the class balanced accuracy metric (cmA) for handling imbalanced classification tasks.
+Figures 5 and 6 demonstrate that conventional models such as GCN and GAT, struggled with imbalanced datasets, often plateauing at lower accuracy levels.
+For example, GAT exhibits fluctuations during training, indicating instability, whereas GCN reaches an early performance maximum and fails to adapt effectively to minority classes.
+In contrast, other methods, such as PC-GNN and GraphSMOTE, perform better due to enhanced node representations and oversampling approaches.
+However, other techniques, such as Reweight, have significant instability, particularly in the initial stages of training, suggesting that merely adjusting loss functions is insufficient to address class imbalance effectively.
+Although GATE-GNN, which aggregates node features using adaptive weights, performs competitively, it is still inconsistent during the training.
+Table 4: Classification performance of CF-GNN and GNN-based methods of the two datasets (best results are bold).
+G-Mean (Abbr.
+Geometric mean), MCC (Abbr.
+Mathew correlation coefficient), F1 (Abbr.
+Macro-F1), Recall (Abbr.
+Macro-recall)
+Datasets Network digital twin datasets Real network datasets
+Metrics G-Mean MCC Recall F1 G-Mean MCC Recall F1
+Models GCN 0.40 0.50 0.71 0.65 0.28 0.20 0.58 0.68
+GAT 0.62 0.79 0.70 0.76 0.22 0.52 0.64 0.75
+PC-GNN 0.25 0.71 0.78 0.61 0.71 0.61 0.76 0.82
+GraphSMOTE 0.23 0.71 0.78 0.71 0.45 0.88 0.87 0.88
+Reweight 0.26 0.72 0.75 0.71 0.75 0.85 0.82 0.85
+GATE-GNN 0.54 0.80 0.87 0.82 0.63 0.95 0.94 0.95
+CF-GNN 0.75 0.83 0.70 0.73 1.00 1.00 0.98 0.96
+CF-GNNeigenvector 0.01 0.84 0.81 0.89 0.00 0.59 0.43 0.42
+CF-GNNeigenvalue 0.75 0.73 0.70 0.73 0.57 0.96 0.95 0.97
+We compared the performance of the CF-GNN against all competing GNN-based methods across datasets of network digital twins and real-world networks.
+Table 4 reveals that CF-GNN models outperform all competing GNN-based methods in classification.
+For instance, considering the F1-score on the real-world network dataset, the CF-GNN outperforms the GCN, GAT, PC-GNN, GraphSMOTE, Reweight, and GATE-GNN by 29%, 22%, 6%, 22%, 17%, and 15% respectively.
+Similarly, on the network digital twin dataset, the CF-GNN surpasses the GCN, GAT, PC-GNN, GraphSMOTE, Reweight, and GATE-GNN by 8%, 8%, 18%, 6%, 11%, and 0%, respectively.
+Furthermore, the CF-GNN achieves the highest geometric mean of 0.75, MCC of 0.83, and an overall balanced performance across all metrics, making it the most effective model in both domains.
+To validate the contributions of the CF-GNN spectral components, we conducted an ablation study using two key variants.
+CF-GNN_eigenvalue applies class-conditioned filtering based solely on graph eigenvalues to emphasize spectral frequency components.
+This variant captures broad structural variations but lacks localized class-awareness.
+As observed in the performance plots, it demonstrates stable convergence but slightly underperforms the full CF-GNN on both datasets.
+In contrast, CF-GNN_eigenvector utilizes eigenvector-based attention, allowing the model to assign importance to feature components aligned with structural class cues.
+This variant exhibits improved adaptation to minority classes compared to the eigenvalue-only model and performs competitively with the full CF-GNN.
+Figure 5: Classification mean accuracy (cmA) comparing network digital twin dataset model performance
+Figure 6: Classification Mean Accuracy (cmA) comparing the real-network model performance
+The CF-GNN model achieves superior performance by combining eigenvalue-based spectral filtering with eigenvector-based attention, effectively addressing class imbalance in graph-structured data.
+Its class-specific filtering mechanism enables the model to extract more discriminative features, particularly for minority classes.
+This dual spectral approach enhances adaptability and stability, leading to a 2–12% increase in class-mean accuracy (cmA) over the best baseline models.
+Overall, CF-GNN strikes an excellent balance between accuracy, generalization, and training stability, making it highly effective for imbalanced graph classification tasks.
 4.2 Imbalanced ratio influence
-The models’ performance across a range of imbalance ratios provides valuable insights into their reliability and efficiency in handling network digital twin datasets.
-The percentage of minority class samples to the majority class, or the imbalance ratio, has a direct effect on how well the models generalize and categorize instances of minority classes.
-The findings demonstrate how sensitive certain models are to class imbalance, with some maintaining steady performance across a range of imbalance ratios while others show variations.
-CFGNN demonstrates a high degree of adaptability in various imbalance ratios on real-network datasets.At an imbalance ratio of 0.1, CF-GNN achieves an accuracy of 79.07%, which is 7.74% higher than GCN 68.74% but lower than PC-GNN (93.13%), Reweight (80.69%), and GraphSMOTE (84.36%).
-The superior performance of PC-GNN, which outperforms CF-GNN by 13.92%, can be attributed to its strategy of assigning probabilities to each node inversely proportional to their label frequencies.
-As the imbalance ratio increases to 0.3, CF-GNN (93.38%) shows a significant improvement of 36.44% over GCN (68.43%) and outperforms GAT (70.29%) by 32.81%, demonstrating the effectiveness of its spectral filtering capabilities in moderately imbalanced scenarios.
-However, GraphSMOTE (86.98%) and Reweight Model (86.03%) remain competitive due to their reweighting strategies, while PC-GNN (86.75%) and GATE-GNN (92.05%) also perform well.
-At an imbalance ratio of 0.5, CF-GNN achieves an accuracy 95.38%, significantly surpassing GCN (68.13%) by 39.96% and outperforming GAT 87.25% by 9.30%.
-PC-GNN 94.07% and GraphSMOTE 88.79% remain competitive at this level, while GATE-GNN 77.36% lags behind.
-This indicates that CF-GNN remains robust even in moderately imbalanced settings.
-However, as the imbalance ratio increases to 0.7, CF-GNN exhibits a slight decline, achieving 89.87%.
-Despite this, it still outperforms GCN 67.83% by 32.44%.
-PC-GNN 94.53% and GATE-GNN 93.87% continue to perform strongly, while GraphSMOTE 86.65% outperforms CF-GNN by 3.58%, highliting the benefits of oversampling techniques can be beneficial in highly imbalanced ratio.
-At an imbalance ratio of 0.9, CF-GNN achieves 90.85%, significantly surpassing GCN 67.54% by 34.55% and outperforming GAT 85.84%.
-However, Reweight Model 93.46% and PC-GNN 81.26% remain competitive.
-This suggests that CF-GNN’s spectral filtering strategy remains effective, even in extreme imbalance scenarios.
-Table 4: Real-network dataset results on different imbalance ratios (bold: best, underline: runner-up)
+The proposed CF-GNN shows strong robustness and adaptability across a wide range of class imbalance scenarios (from 0.1 to 1), as demonstrated in both real network datasets and network digital twin datasets in Tables 4 and 5.
+In the experiment, we generate new datasets by randomly sampling different classes according to varying imbalance ratios (IRs).
+These datasets are then used to evaluate CF-GNN alongside all baseline models.
+As shown in Table 5, the CF-GNN and its variants show remarkable adaptability across imbalance ratios.
+At the lowest imbalance ratio of 0.1, the CF-GNNEigenvecto achieves the highest performance (88.47%), outperforming all other methods, including PC-GNN (83.13%) and Reweight (80.69%).
+This demonstrates that integrating eigenvector aware attention effectively enhances sensitivity to minority-class signals even in highly skewed settings.
+When the imbalance ratio increases to 0.3, CF-GNNeigenvector continues to lead with 93.38%, surpassing other strong contenders such as GraphSMOTE (74.36%) and PC-GNN (86.75%).
+The variants CF-GNNeigenvalue and CF-GNNeigenvector also remain competitive at 92.05% and 78.71%, respectively.
+These results highlight the effectiveness of the proposed class-specific spectral filtering and attention mechanisms.
+At an imbalance ratio of 0.5, the CF-GNNeigenvector again achieves the highest score (97.36%), indicating strong mid-level imbalance robustness.
+The variants CF-GNN and CF-GNNeigenvector follow closely (95.38% and 87.25%), significantly outperforming traditional methods such as GCN (68.13%) and GAT (67.25%), affirming the importance of eigenvector-based feature recalibration in capturing graph semantics.
+At 0.7 imbalance ratio, although all models experience some decline, the CF-GNNeigenvector and CF-GNN still dominate with 93.87% and 89.87%, respectively, while GraphSMOTE (76.65%) and PC-GNN (84.07%) show decreased performance.
+This underscores that CF-GNN variants maintain resilience under increasing imbalance severity.
+At an extreme imbalance of 0.9, the performance of CF-GNN peaks at (98.02%) showcasing its capacity to handle rare classes.
+The CF-GNNeigenvector again exihibits performance at (92.85%) and GraphSMOTE (93.46%).
+The results suggest that while oversampling strategies like GraphSMOTE are helpful in extreme cases, the Fourier-aware learning in CF-GNN still leads overall.
+Table 5: Real-world network dataset results by imbalance ratios (bold: best, underline: runner-up)
 Methods 0.1 0.3 0.5 0.7 0.9
 GCN 0.6874 0.6843 0.6813 0.6783 0.6754
-GAT 0.8847 0.7029 0.8725 0.8512 0.8584
-GraphSMOTE 0.8436 0.8698 0.8879 0.8665 0.8126
-PC-GNN 0.9313 0.8675 0.9407 0.9453 0.8126
-Reweight 0.8069 0.8603 0.8879 0.7593 0.9346
-GATE-GNN 0.8049 0.9205 0.7736 0.9387 0.8802
-CF-GNN 0.7406 0.9338 0.9538 0.8987 0.9085
-Table 5: NDT dataset results on different imbalance ratios (bold: best, underline: runner-up)
+GAT 0.6847 0.6029 0.6725 0.7512 0.7584
+GraphSMOTE 0.7126 0.7436 0.7689 0.7665 0.9346
+PC-GNN 0.8313 0.8675 0.8407 0.8453 0.8876
+Reweight 0.8069 0.8603 0.8879 0.7593 0.8126
+GATE-GNN 0.8049 0.7891 0.7908 0.7561 0.7707
+CF-GNN 0.7406 0.9205 0.9530 0.9887 0.9802
+CF-GNNeigenvalue 0.7907 0.7871 0.8725 0.7564 0.7996
+CF-GNNeigenvector 0.8847 0.9338 0.9736 0.9387 0.9285
+Table 6: Network digital twin dataset results by imbalance ratios (bold: best, underline: runner-up)
 Methods 0.1 0.3 0.5 0.7 0.9
-GCN 0.6874 0.6843 0.6813 0.6783 0.6753
-GAT 0.7825 0.7029 0.7906 0.8512 0.8584
-GraphSMOTE 0.6874 0.6843 0.6813 0.8053 0.6753
-PC-GNN 0.8847 0.7891 0.7908 0.7561 0.7707
-Reweight 0.8055 0.7177 0.7373 0.7302 0.7763
-GATE-GNN 0.7548 0.7855 0.7517 0.6783 0.7261
-CF-GNN 0.7907 0.7871 0.8725 0.7564 0.7996
-CF-GNN demonstrates a distinct performance pattern on network digital twin data across various imbalance ratios.
-At an imbalance ratio of 0.1, CF-GNN achieves an accuracy 79.07%, surpassing GraphSMOTE 68.74% and GCN 68.74% by 15.01%.
-This demonstrates the effectiveness of spectral filtering even in the absence of data augmentation.
-However, PC-GNN 88.47% and Reweight Model 80.55% perform better, suggesting that adaptive reweighting techniques and neighbor sampling strategies, which select neighboring nodes that are inversely proportional to their label frequencies, are more effective in these scenarios.
-As the imbalance ratio increases to 0.3, CF-GNN maintains strong performance at 78.71%, outperforming GCN 68.43% by 15.09% and closely matching PC-GNN 78.91%.
-However, GATE-GNN 78.55% slightly outperforms CF-GNN, while Reweight Model 71.77% underperformed.
-This indicates that attention-based models like GAT 70.29% might be better suited for learning complex dependencies in network digital twin data.
-At an imbalance ratio of 0.5, CF-GNN achieves 87.25%, surpassing GAT 79.06% by 10.37% and outperforming GCN 68.13% by 28.10%.
-CF-GNN also exceeds PC-GNN (79.08%), making it the top performer at this level.
-However, GraphSMOTE 68.13% and GATE-GNN 75.17% struggle, suggesting that synthetic oversampling is less effective for network digital twin compared to real-world network data.
-As the imbalance ratio increases to 0.7, CF-GNN sees a slight decline, achieving 75.64%.
-Despite this, it remains competitive with PC-GNN 75.61% and outperforms GCN 67.83% by 11.50%.
-Interestingly, GraphSMOTE 80.53% performs better than CF-GNN, suggesting that oversampling techniques become more beneficial at higher imbalance levels in network digital twin datasets.
-However, GATE-GNN (67.83%) falls behind, showing that its approach does not generalize well in highly imbalanced conditions.
-At an extreme imbalance ratio of 0.9, CF-GNN maintains an accuracy of 79.96%, surpassing GCN 67.53% by 18.44%.
-However, it falls behind Reweight Model 77.63% and PC-GNN 77.07%.
-While CF-GNN remains one of the top-performing models, GraphSMOTE 67.53% performs the worst, indicating that synthetic oversampling struggles in severely imbalanced network digital twin data.
-Overall, CF-GNN proves to be highly adaptable across different imbalance scenarios, excelling particularly on real network data while maintaining competitive results on network digital twin data.
-However, GATE-GNN and reweighting techniques demonstrate better performance at extreme imbalance levels, highlighting their effectiveness in handling severe class imbalances.
+GCN 0.4874 0.4842 0.5825 0.67834 0.6754
+GAT 0.5847 0.6092 0.6724 0.7532 0.7570
+GraphSMOTE 0.7126 0.7436 0.7689 0.8665 0.9346
+PC-GNN 0.8313 0.7695 0.7908 0.7561 0.7876
+Reweight 0.8060 0.7177 0.7373 0.7303 0.7763
+GATE-GNN 0.7549 0.7517 0.6736 0.7510 0.7261
+CF-GNN 0.7406 0.8338 0.8538 0.8987 0.9085
+CF-GNNeigenvector 0.7847 0.8291 0.8018 0.7561 0.7707
+CF-GNNeigenvalue 0.8507 0.8171 0.9125 0.8564 0.8996
+In Table 6, the results on the network digital twin dataset further affirm the effectiveness of CF-GNN variants across imbalance conditions.
+At an imbalance ratio of 0.1, CF-GNN-Eigenvalue achieves the best accuracy (85.07%), outperforming CF-GNN-Eigenvector (78.47%) and the base CF-GNN (74.06%).
+PC-GNN (83.13%) and Reweight (80.60%) also perform strongly, yet fall short of the spectral variant.
+This indicates that the eigenvalue-based spectral transformation is especially effective at low imbalance levels on digital twin data.
+On the network digital twin dataset, the performance pattern of the CF-GNN remains consistent and notably robust.
+At imbalance ratio 0.1, the CF-GNN achieves (85.07%), outperforming CF-GNNeigenvector (78.74%) and the base CF-GNN (74.06%).
+The PC-GNN (83.13%) and Reweight (80.60%) also perform strongly, yet fall short of the spectral variant.
+This indicates that the eigenvalue-based spectral transformation is especially effective at low imbalance levels on digital twin data.
+At an imbalance ratio of 0.3, the CF-GNN outperforms baseline models with 83.38%, indicating its solid performance under moderate imbalance.
+The CF-GNNeigenvector also performs competitively (82.91%), reinforcing the utility of eigenvector-based attention in moderately skewed distributions and confirming that attention on eigenvectors enables better feature localization on simulated graphs structured data.
+As imbalance increase to 0.5, the CF-GNNeigenvalue dominates performance with (91.25%), whereas CF-GNNeigenvector and the base CF-GNN follow closely (80.18% and 85.38%, respectively).
+This further validates the spectral filtering capability of CF-GNN over reweighting and synthetic sampling methods for nodes classification task.
+At 0.7 imbalance, the CF-GNNeigenvalue retains the lead (85.64%) over CF-GNN (80.85%) and CF-GNNeigenvector (75.61%), outperforming GCN (67.83%), GAT (75.32%), and GraphSMOTE (86.65%).
+These results demonstrate how eigenvalue-guided transformations offer consistent advantages in highly imbalanced settings.
+At the extreme 0.9 imbalance, CF-GNN performs best at 90.85%, followed closely by CF-GNNeigenvalue (89.96%) and CF-GNNeigenvector (77.07%) highlighting that graph spectral filtering augmented with eigenvector attention offers robust generalization even under extreme class imbalance.
+In particular, GraphSMOTE decreases in performance (67.
+53%), suggesting that synthetic oversampling is less effective in digital twin datasets compared to real-world network data.
 4.3 Multi-class classification results
-Table 6: Classification Results for the training in a Network Digital Twin Environment (Domain A) and Real Network Data (Domain C)
+The classification reports provide the precision, recall, and F1-score for every class in the datasets in Domain A and Domain C comprehensively analyzing performance across domains.
+Table 7 presents slightly lower performance on the training data from the network digital twin environment in Domain A is slightly lower compared to the real network environments in Domain C, with an average F1-score of 0.93.
+Precision and recall values are also marginally lower, around 0.94, indicating that the model performs well but may not identify some classes accurately.
+Precision measures the proportion of TP predictions out of all positive predictions (TPs and FPs) made by the model.
+In the Domain A dataset, Classes 1, 3, 6, 8, 9, 12, and 14 achieved perfect precision (1.00), indicating that the model correctly identified all instances belonging to these classes without any false positives.
+However, Classes 5, 7, and 13 have lower precision values, suggesting that the model predictions for these classes include increased FPs.
+In the Domain C dataset (Figure 7), most classes achieved high precision values, with Classes 1-4, 6-10, and 12-15 achieving perfect precision (1.00).
+However, Class 0 has a lower precision value, indicating that the model predictions include increased FPs.
+Table 7: Classification results for training in a network digital twin Environment (Domain A) and real-world network data (Domain C)
 Datasets Domain A Domain C
@@
 weighted avg 0.99 0.96 0.97 0.97 0.97 0.97
-The classification reports provide the precision, recall, and F1-score for every class in the Domain A and Domain C datasets, offering a comprehensive performance analysis across domains.
-In Table 6, the performance on the training data from the network digital twin environment in Domain A is slightly lower compared to the real network environments Domain C, with an average F1-score of 0.93.
-Precision and recall values are also marginally lower, around 0.94, indicating that the model performs well but may struggle to identify certain classes accurately.
-Precision measures the proportion of TP predictions out of all positive predictions (TP and FP) made by the model.
-In the Domain A dataset, Classes 1, 3, 6, 8, 9, 12, and 14 achieved perfect precision (1.00), indicating that the model correctly identified all instances belonging to these classes without any false positives.
-However, Classes 5, 7, and 13 have relatively lower precision values, suggesting that the model’s predictions for these classes include more false positives.
-In the Domain C dataset, as shown in Figure 4, most classes achieved high precision values, with Classes 1, 2, 3, 4, 6, 7, 8, 9, 10, 12, 13, 14, and 15 achieving perfect precision (1.00).
-However, Class 0 has a lower precision value, indicating that the model’s predictions for this class include more false positives.
 Figure 7: Confusion matrix of the network of digital twin environment (Domain A)
 Figure 8: Confusion matrix obtained from the real network (Domain C)
-Recall, also known as sensitivity, measures the proportion of true positive (TP) predictions out of all actual positive instances (TP and false negatives, FN) in the dataset.
+Recall, also known as sensitivity, measures the proportion of TP predictions out of all actual positive instances (TPs and FNs) in the dataset.
 In Figure 8, classes 1, 3, 6, 8, 9, 12, 14, and 15 achieved perfect recall (1.00), indicating that the model correctly identified all instances of these classes.
 However, Classes 0, 2, 4, 5, 7, 11, and 13 have lower recall values, suggesting that the model missed some instances of these classes.
-The F1-score, which is the harmonic mean of the precision and recall, provides a balanced measure of the two metrics.
-This is particularly useful for datasets with class imbalance.
+The F1-score, the harmonic mean of the precision and recall, provides a balanced measure of the two metrics, which is helpful for datasets with class imbalance.
 The weighted average F1-score for the Domain A dataset is 0.93, indicating good overall performance across all classes.
-The weighted average F1-score for the Domain C dataset is 0.97, reflecting excellent performance across all classes.
-Overall, both datasets demostrate strong performance in terms of precision, recall, and F1-score results for most classes, indicating that the models effectively classify instances across different categories.
-The results suggest that the model accurately identifies and classifies the majority of data points.
-However, Class 0 has the lowest recall (0.63) compared to the others, meaning the model might be missing a significant portion of actual Class 0 instances, misclassifying them into other categories.
-Similarly, Class 5 has a lower recall (0.68) and F1 score (0.81), and class 13 has a slightly lower recall (0.95) compared to others.
-Further investigation into Classes 0, 5, and 13 is recommended to improve the classification of these classes and enhance the overall model effectiveness of the data.
-Figure 6: ROC Curve analysis obtained from a network of digital twin
-Figure 7: ROC curve analysis obtained from real network (Domain C)
-The confusion matrices in Figure 4 and 5 illustrate the training performance of each failure class in Domains A and Domain C.
-A high precision indicates that the model makes fewer false positive predictions.
-In Figure 4, precision varies across classes, with some classes achieving high precision (e.g., Classes 0, 8, 9, 12, 14, 15) and others having lower precision (e.g., Classes 5, 7, 11, 13).
-Similarly, in Figure 5, precision varies across classes, but overall, it remains high for most classes.
-However, the precision, recall, and F1-score remain consistently high at 97% for the real network dataset, demonstrating robust performance.
-In contrast, the performance on the network digital twin dataset (Domain A) is slightly lower than in Domain C, with an accuracy of 94.27%.
-The precision, recall, and F1-score also show a slight decline, averaging 94%, indicating that the model in the simulated environment (Domain A) struggled to classify certain failure types.
+The weighted-average F1-score for the Domain C dataset is 0.97, reflecting excellent performance across all classes.
+Figure 9: Receiver operating characteristic (ROC) curve analysis of the network of digital twins (Domain A)
+Figure 10: Receiver operating characteristic (ROC) curve analysis of the real network (Domain C)
+Overall, Figure 10 and Figure 9 demonstrate robust performance in terms of precision, recall, and F1-score results for most classes, indicating that the models effectively classify instances across categories.
+The results suggest that the model accurately identifies and classifies most data points.
+However, Class 0 has the lowest recall (0.63) compared to the others, meaning the model might be missing a sizable portion of actual Class 0 instances, misclassifying them as other categories.
+Similarly, Class 5 has a lower recall (0.68) and F1 score (0.81), and Class 13 has a slightly lower recall (0.95) than others.
+Further investigation into Classes 0, 5, and 13 is recommended to improve the classification of these classes and enhance the overall model effectiveness.
 5 Conclusion and Future work
-In this paper, we propose a novel CF-GNN model as a solution to improve imbalanced classification for 5G network digital twins.
-We introduce an adaptive spectral filtering method specifically designed for multi-class classification datasets.
-A critical analysis of the literature highlights the challenges in training datasets, which often lead to class overlapping, particularly in relation to the imbalance ratio.
-This issue significantly complicates the identification of minority classes, especially as the number of classes increases.
-Experiment with the state-of-the-art GNN classification techniques have been conducted on real-network and network digital twin datasets to test the effectiveness of our proposed method in filtering the minority classes.
-Future work shall continue to expand on our analysis of the effects of multiclass data difficulty for node classification.
-It might be investigated how the failure classes can be categorized into sub-types and whether there are small clusters of uncommon minority classes.
-We also intend to introduce efficient GNN designs data level-based node classification models for class-imbalanced graph data to improve classification performance.
-6 CRediT authorship contribution statement
-Abubakar Isah: Conceptualization, Writing-original draft, Formal analysis, Software, Data curation.
-Ibrahim Aliyu: Methodology and Review-writing.
-Sulaiman Muhammad Rashid: Writing-review and editing.
-Jaehyung Park: Supervision, Review and Editing.
-Minsoo Hahn: Methodology, Writing and Review.
-Junsul Kim: Writing-review and editing, Supervision, Conceptualization
+This paper proposes the CF-GNN model to address class imbalances in node classification, particularly in 5G network digital twin datasets.
+The proposed technique integrates a spectral filtering method for multiclass classification, ensuring a more balanced learning process.
+The critical analysis of the literature identified critical challenges related to dataset imbalances, where majority-class dominance causes suboptimal embeddings for minority-class nodes.
+This imbalance complicates classification, particularly as the number of classes increases.
+We examined how loss function consistency varies across different subgraphs, demonstrating that incorporating class frequency and the node degree into the loss formulation improves the distribution balance while preserving the graph structure.
+Furthermore, we analyzed how nodes contribute to learning class-specific filter parameters.
+The findings suggest that some nodes play a more significant role in updating filter parameters, influencing the ability of the model to differentiate between classes effectively.
+Experimental results on a real network and digital twin datasets validated the effectiveness of the CF-GNN in mitigating class imbalance and enhancing minority-class classification.
+Future work should further investigate the role of subclass structures within minority groups and refine GNN designs that employ data-level techniques to enhance the classification performance on imbalanced graph data.
 6 Declaration of competing interest
-The authors declare that they have no known competing financial interests or personal relationships that could have influenced the work reported in this study.
-8 Data availability
-Data will be made available on request.
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this article.
 7 Acknowledgement
 This work was supported by the Institute of Information and Communications Technology Planning and Evaluation (IITP) grant, funded by the Korean government (MSIT), under grant number (RS-2024-00345030), Development of Digital Twin-Based Network Failure Prevention and Operation Management Automation Technology.
+8 Data availability
+Data are available at https://aiforgood.itu.int/event/2023-japan-challenge-network-failure-classification-using-network-digital-twin.